### v1

# Hard View Selection for Contrastive Learning

Abstract
Many Contrastive Learning (CL) methods train their models to be invariant to different “views” of an image input for which a good data augmentation pipeline is crucial. While considerable efforts were directed towards improving pre-text tasks, architectures, or robustness (e.g., Siamese networks or teacher-softmax centering), the majority of these methods remain strongly reliant on the random sampling of operations within the image augmentation pipeline, such as the random resized crop or color distortion operation. In this paper, we argue that the role of the view generation and its effect on performance has so far received insufficient attention. To address this, we propose an easy, learning-free, yet powerful Hard View Selection (HVS) strategy designed to extend the random view generation to expose the pretrained model to harder samples during CL training. It encompasses the following iterative steps: 1) randomly sample multiple views and create pairs of two views, 2) run forward passes for each view pair on the currently trained model, 3) adversarially select the pair yielding the worst loss, and 4) run the backward pass with the selected pair. In our empirical analysis we show that under the hood, HVS increases task difficulty by controlling the Intersection over Union of views during pretraining. With only 300-epoch pretraining, HVS is able to closely rival the 800-epoch DINO baseline which remains very favorable even when factoring in the slowdown induced by the additional forwards of HVS. Additionally, HVS consistently achieves accuracy improvements on ImageNet between 0.55% and 1.9% on linear evaluation and similar improvements on transfer tasks across multiple CL methods, such as DINO, SimSiam, and SimCLR.
1 Introduction
Figure 1: HVS first samples $N$ views, pairs them, and adversarially selects the hardest pair, i.e., the one with the worst loss according to the current model state.
Various approaches to learn effective and generalizable visual representations in Self-Supervised Learning (SSL) exist. One way to categorize many SSL methods is to distinguish generative and discriminative approaches. While generative methods aim at generating image input, discriminative methods, with contrastive learning (Hadsell et al., 2006; He et al., 2020) being the most renowned approach, aim at learning a latent representation in which similar image views are closely and dissimilar ones are distantly located.
Such views are generated by applying a sequence of (randomly sampled) image transformations and are usually composed of geometric (cropping, rotation, etc.) and appearance (color distortion, blurring, etc.) transformations. A body of literature (Chen et al., 2020a; Wu et al., 2020; Purushwalkam & Gupta, 2020; Wagner et al., 2022; Tian et al., 2020b) has illuminated the effects of image views on representation learning and identified random resized crop (RRC) transformation, which randomly crops the image and resizes it back to a fixed size, as well as color distortion as one of the most essential ones for effective contrastive learning. However, despite this finding and to our best knowledge, little research has gone into identifying more effective ways for selecting or generating views to improve performance. In this paper, we seek to extend the random view generation to expose the pretrained model to harder samples during CL training.
Figure 2: Example images (left) and sampled views (right) after resizing and appearance transformations. Pairs selected adversarially by HVS are shown with a solid frame.
Supervised Learning works that address this shortcoming in a distant manner and which attempt to leverage task difficulty include Spatial Transformer Networks (Jaderberg et al., 2015)and Adversarial AutoAugment (Zhang et al., 2020), which either directly learn geometric transformations or generate augmentation policies that yield hard samples for more robust feature learning. However, these approaches have not been used successfully for unsupervised learning and Adversarial AutoAugment is limited in leveraging the geometric perspective for view hardness control as it does not include RRC in its search space. Furthermore, and in contrast to what we propose, integrating any of these methods into existing contrastive learning pipelines is non-trivial.
Contrastive learning approaches that are related to our work and aim at controlling view hardness include (Shi et al., 2022) and (Tamkin et al., 2021) which propose masking or perturbing input images through adversarial models. Another approach is MADAug (Hou et al., 2023) which use a cooperative setting to learn policies yielding progressively harder samples based on a validation loss which is usually unavailable in contrastive learning. A work that takes a similar perspective like us in leveraging the information content of views for contrastive learning is (Tian et al., 2020b). The authors unify the data augmentation of several CL methods under the lens of mutual information and learn view generators under this premise. While these works offer important insights into view optimality, they usually require learning additional models or non-trivial changes to the pipeline and thus, are not easily integrated into existing CL methods.
In contrast, we propose a fully learning-free, easy-to-integrate approach that applies the following simple strategy: we sample a few views and forward each pair through the currently trained model. Next, we identify the pair that yields the worst loss among all pairs, use it for the backward pass, and repeat (see Fig. 1). While being agnostic to any contrastive loss, this strategy exposes the pretrained model to samples that the model at any given point in the training trajectory finds challenging. As our results show, sampling only four views are sufficient to achieve noticeable improvements.
Overall, our contributions can be summarized as follows:
  • We propose Hard View Selection (HVS), a new learning method complementing contrastive learning that is easy to integrate and extends the broadly-used random view generation to automatically expose the model to harder samples during pretraining. The only requirement of HVS is that it needs to be able to compute sample-wise losses;
  • We demonstrate the effectiveness and compatibility of our approach using ImageNet (Deng et al., 2009) pretraining across popular and widely-adopted contrastive learning methods that cover a diverse range of objectives such as SimSiam (Chen & He, 2021), DINO (Caron et al., 2021), SimCLR (Chen et al., 2020a);
  • HVS improves all baselines noticeably compared to using the common view generation and we show this on a diverse set of transfer datasets and tasks, including linear evaluation, finetuning, object detection and segmentation;
  • We present insights into the underlying mechanisms of HVS. One key observation is that harder samples are achieved by scheduling the Intersection over Union conditioned on the current model state and image input. We also provide further results on the effects of changing the number of sampled views, color augmentation and studies on longer training.
We make our PyTorch (Paszke et al., 2019) code, models, and all used hyperparameters publicly available under https://anonymous.4open.science/r/hard-view-selection/.
2 Related Work
2.1 Contrastive Learning in SSL
The core idea behind the contrastive learning framework is to learn image representations by contrasting positive pairs (two views of the same image) against negative pairs (two views of different images) (Hadsell et al., 2006). To work well in practice and to prevent model collapse, contrastive learning methods often require a large number of negative samples (Wu et al., 2018; van den Oord et al., 2018; Chen et al., 2020a; He et al., 2020; Tian et al., 2020a; Chen et al., 2020b) stored in memory banks (Wu et al., 2018; He et al., 2020) or, for instance, in the case of SimCLR, implicitly in large batches (Chen et al., 2020a). Non-contrastive approaches, such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), DINO (Caron et al., 2021) and others (Zbontar et al., 2021; Caron et al., 2020; Ermolov et al., 2021), are able to only use positive pairs without causing model collapse but rely on other techniques, such as Siamese architectures, whitening of embeddings, clustering, maximizing the entropy of the embeddings, momentum encoders and more.
2.2 Optimizing for Hard Augmentations and Views in SSL
Past work has shown that optimizing augmentation policies or views directly can be beneficial. Hence, the realm of learning task-specific augmentation policies based on data has seen quick development (Cubuk et al., 2019; Ho et al., 2019; Lin et al., 2019; Zhang et al., 2020; Hataya et al., 2020; Müller & Hutter, 2021). Most related to our work among these approaches is Adversarial AutoAugment (AAA) (Zhang et al., 2020) that proposes to generate augmentation policies yielding hard samples by learning with an adversarial supervised objective. In a similar, but cooperative setting, MADAug (Hou et al., 2023) learns policies yielding progressively harder samples based on a validation loss. In contrast to our work, however, AAA and MADAug do not include the random resize crop operation into their search spaces and thus, are limited in leveraging the geometric perspective to control view hardness.
Unlike learning hard augmentations, approaches exist that optimize directly on the pixel level to control view hardness. For instance, Shi et al. (2022) adversarially learn semantically meaningful masks for a pretrained network to inpaint. Another approach perturbs input images using an adversarial model (Tamkin et al., 2021). Tian et al. (2020b) leverage, like us, the view content for better performance. Under the lens of mutual information (MI), they unify several data augmentation CL methods and derive an MI-based objective to learn view generators for unsupervised and semi-supervised learning. While these methods can complement our approach, they are less easily integrated into existing contrastive pipelines due to the need for adversarial network training.
3 Method
3.1 Contrastive Learning Framework
In this section, we introduce our approach which is also depicted in Algorithm 12. Many different contrastive learning (He et al., 2020) objectives exist, each characterized by variations stemming from design choices, such as by the use of positive and negative samples or asymmetry in the encoder-projector network structure. For simplicity of exposure, we will introduce our approach based on the SimSiam objective Chen & He (2021), but we do note that our method can be used with any other contrastive learning objective that allows the computation of a sample-wise losses.
SimSiam works as follows. Assume a given set of images $\mathcal{D}$, an image augmentation distribution $\mathcal{T}$, a minibatch of $M$ images $\mathbf{x}=\left\{x_{i}\right\}_{i=1}^{M}$ sampled uniformly from $\mathcal{D}$, and two sets of randomly sampled image augmentations $A$ and $B$ sampled from $\mathcal{T}$. SimSiam applies $A$ and $B$ to each image in $\mathbf{x}$ resulting in $\mathbf{x}^{A}$ and $\mathbf{x}^{B}$. Both augmented sets of views are subsequently projected into an embedding space with $\mathbf{z}^{A}=g_{\theta}(f_{\theta}(\mathbf{x}^{A}))$ and $\mathbf{h}^{B}=f_{\theta}(\mathbf{x}^{B})$ where $f_{\theta}$ represents an encoder (or backbone) and $g_{\theta}$ a projector network. SimSiam then minimizes the following objective:
$\mathcal{L}(\theta)=\frac{1}{2}\left(D(\mathbf{z}^{A},\mathbf{h}^{B})+D(\mathbf{z}^{B},\mathbf{h}^{A})\right)$ (1)
where D denotes the negative cosine similarity function. Intuitively, when optimizing $\theta$, the embeddings of the two augmented views are attracted to each other.
3.2 Hard View Selection
As previously described, Hard View Selection extends the random view generation by sampling adversarially harder views during pretraining. Instead of having two sets of augmentations $A$ and $B$, we now sample $N$ sets of augmentations, denoted as $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{N}\}$. Each set $A_{i}$ is sampled from the image augmentation distribution $\mathcal{T}$, and applied to each image in $\mathbf{x}$, resulting in $N$ augmented sets of views $\mathbf{x}^{A_{1}},\mathbf{x}^{A_{2}},\ldots,\mathbf{x}^{A_{N}}$. Similarly, we obtain $N$ sets of embeddings $\mathbf{z}^{A_{1}},\mathbf{z}^{A_{2}},\ldots,\mathbf{z}^{A_{N}}$ and predictions $\mathbf{h}^{A_{1}},\mathbf{h}^{A_{2}},\ldots,\mathbf{h}^{A_{N}}$ through the encoder and projector networks. We then define a new objective function that seeks to find the pair $(x_{i}^{A_{k}},x_{i}^{A_{l}})$ of a given image $x_{i}$ that yields the highest loss:
$(x_{i}^{A_{k}},x_{i}^{A_{l}})=\operatorname*{arg\,max}_{k,l;k\neq l}\mathcal{L}(\theta)_{i,k,l}=\operatorname*{arg\,max}_{k,l;k\neq l}\frac{1}{2}\left(D(z_{i}^{A_{k}},h_{i}^{A_{l}})+D(z_{i}^{A_{l}},h_{i}^{A_{k}})\right),$ (2)
where $\mathcal{L}(\theta)_{i,k,l}$ simply denotes a sample-wise variant of Eq. 1. Overall, we first generate $N$ augmented views for each image $x_{i}$ in the minibatch. Then, we create all combinatorially possible $\binom{N}{2}$ pairs of augmented images for $x_{i}$. Then, we use Eq. 2 to compute the sample-wise loss for each pair which corresponds to invoking forward passes. We then select all pairs that yielded the highest loss to form the new hard minibatch of augmented sets $\textbf{x}^{A_{k_{*}}}$ and $\textbf{x}^{B_{l_{*}}}$, discard the other pairs and use the hard minibatch for optimization. As shown in Algorithm 12, we repeat this process in each training iteration. This modification introduces a more challenging learning scenario in which the model is encouraged to learn more discriminative features by being exposed to harder views. While the additional forward passes come with a certain overhead (about a factor of 1.55$\times$ for SimSiam), as we will demonstrate in our experiments, the focus on harder views leads to consistent improvements.
1:Input: Number of views $N\geq 2$, batch size $M$,
2:augmentation distribution $\mathcal{T}$, encoder $f_{\theta}$, projector $g_{\theta}$
3:for each $x_{i}$ in the sampled minibatch $\left\{x_{i}\right\}_{i=1}^{M}$ do
4:     Sample $N$ augmentations: $A=\left\{t_{n}\sim\mathcal{T}\right\}_{n=0}^{N}$
5:     Create augmented views: $\textbf{x}_{i}^{A}=\left\{t_{n}(x_{i})\right\}_{n=0}^{N}$
6:     Create all $\binom{N}{2}$ possible pairs $\textbf{x}_{i}^{A_{k}}\times\textbf{x}_{i}^{A_{l}}$, $k\neq l$
7:     For all pairs $(x_{i}^{A_{k}},x_{i}^{A_{l}})$ forward through $f_{\theta}$ and $g_{\theta}$
8:     Select the hard pair $(x_{i}^{A_{k_{*}}},x_{i}^{A_{l_{*}}})$ maximizing Eq. 2
9:Create the minibatch $(\textbf{x}^{A_{k_{*}}},\textbf{x}^{A_{l_{*}}})$ consisting of hard pairs
10:Proceed with standard contrastive learning training
to update $\theta$ with the hard batch
11:Repeat for all minibatches
12:return Optimized $\theta$
Algorithm 1 Contrastive Learning with HVS
While we exemplified the integration of HVS with the SimSiam objective, integrating it into other contrastive methods is as straightforward. The only requirement of HVS is to be able to compute sample-wise losses (in order to select the views with the highest loss). In our experiments section and in addition to SimSiam, we study the application of HVS to the objectives of DINO and SimCLR (see also Appendix H.1 for a formal exemplary definition for the integration of HVS into SimCLR).
3.3 Implementation and Evaluation Protocols
Implementation
We now describe the technical details of our approach. HVS can be used with any contrastive method that allows computing sample-wise losses, and the only two elements in the pipeline we adapt are: 1) the data loader (needs to sample $N$ views for each image) and 2) the forward pass (invokes the select function to identify and return the hard views). The image transformation distribution $\mathcal{T}$ of the baselines is left unchanged. Note, for the view selection one could simply resort to RRC only and apply the rest of operations after the hard view selection (see Section 5.1 for a study on the influence of appearance on the selection). All experiments were conducted with $N=4$ sampled views, yielding $\binom{N}{2}=6$ pairs to compare, except for DINO that uses 10 views (2 global, 8 local heads) per default. For DINO, we apply HVS to both global and local heads but to remain tractable, we upper-bound the number of total pair comparisons to 128. SimCLR uses positive and negative samples. In accordance with the simplicity of HVS, we do not alter its objective, which naturally leads to selecting hard views that are adversarial to positive and “cooperative” to negative views.
Method Arch. 100 epochs 300 epochs
Lin. k-NN Lin. k-NN
DINO ViT-S/16 73.52 68.80 75.48 72.62
+ HVS ViT-S/16 74.67 70.72 76.56 73.64
Improvement +1.15 +1.92 +1.08 +1.02
DINO RN50 71.93 66.28 75.25 69.53
+ HVS RN50 72.87 67.33 75.80 70.11
Improvement +0.94 +1.05 +0.55 +0.58
SimSiam RN50 68.20 57.47 70.35 61.40
+ HVS RN50 68.98 58.97 70.90 62.97
Improvement +0.78 +1.50 +0.55 +1.57
SimCLR RN50 63.40 52.83 65.50 55.65
+ HVS RN50 65.33 54.76 67.20 57.31
Improvement +1.93 +1.85 +1.70 +1.66
Table 1: Average top-1 linear and k-NN classification accuracy on the ImageNet validation set for 100 and 300-epoch pretrainings. For SimCLR we run 200 instead of 300 epochs.
Evaluation Protocols
We now describe the protocols used to evaluate the performance in our main results section. In self-supervised learning, it is common to assess pretraining performance with the linear evaluation protocol by training a linear classifier on top of frozen features or finetuning the features on downstream tasks. Our general procedure is to follow the baseline methods as closely as possible, including hyperparameters and code bases (if reported). It is common to use RRC and horizontal flips during training and report the test accuracy on central crops. Due to the sensitivity of hyperparameters, and as done by Caron et al. (2021), we also report the quality of features with a simple weighted nearest neighbor classifier (k-NN).
4 Main Results
Here, we discuss our main results on image classification, object detection, and segmentation tasks. All results are self-reproduced using the baseline code and available hyperparameters, and are averaged over multiple seeds. All hyperparameters are reported in Appendix G.
4.1 Evaluations on ImageNet
We report the top-1 validation accuracy on frozen features, as well as the k-NN classifier performance, in Table 1. For DINO, we additionally compare ResNet-50 (He et al., 2016) against the ViT-S/16 (Dosovitskiy et al., 2020) architecture. Our method compares favourably against all baselines with an increased performance of approximately $1\%$ on average, showing the benefit of sampling hard views. We highlight that our 300-epoch DINO+HVS model exhibits strong performance, closely rivaling the officially reported 800-epoch DINO performance of 77%, with a mere ${\sim}0.4\%$ performance gap. This achievement remains favorable, even when factoring in the 2$\times$ slowdown (DINO with 2 global, 8 local heads) induced by our method. Due to limited compute resources, we run the pretrainings for 100 and 300 epochs (200 epochs for SimCLR) and batch sizes of 512 (100 epoch) or 1024 (200 & 300 epoch trainings), respectively. This choice is in line with a strategy that favors the evaluation of a diverse and larger set of baselines over the evaluation of a less diverse and smaller set and underpins the broad applicability of HVS. We primarily ran our experiments with 8xNVIDIA GeForce RTX 2080 Ti nodes, with which the cheapest runs required ${\sim}3.5$ days for pretraining and linear evaluation and the most expensive runs required ${\sim}25$ days. Moreover, we emphasize that HVS is insensitive to the baseline hyperparameters and simply reusing these consistently resulted in improvements of the reported magnitudes across all experiments.
4.2 Transfer to Other Datasets and Tasks
We now demonstrate the transferability of features learned with HVS. For all transfer experiments, we use our 100-epoch ImageNet pretrained SimSiam ResNet-50 and DINO ViT-S/16 models.
Method Arch. CIFAR10 CIFAR100 Flowers102 iNat 21 Food101
Lin. F.T. Lin. F.T. Lin. F.T. Lin. F.T. Lin. F.T.
SimSiam RN50 82.60 95.50 54.20 77.20 34.27 56.40 32.50 60.30 65.70 83.90
+ HVS RN50 84.40 96.10 57.10 78.20 38.37 58.90 33.90 60.90 67.10 84.70
Improvement +1.80 +0.60 +2.90 +1.00 +4.10 +2.50 +1.40 +0.60 +1.40 +0.80
DINO ViT-S/16 94.53 98.53 80.63 87.90 91.10 93.20 46.93 53.97 73.30 87.50
+ HVS ViT-S/16 95.13 98.65 81.27 88.23 92.07 93.60 49.03 54.16 74.13 87.91
Improvement +0.60 +0.12 +0.63 +0.33 +0.97 +0.40 +2.10 +0.19 +0.83 +0.41
Table 2: Models trained on ImageNet with HVS compare favourably against models trained without it when transferred to other datasets through finetuning (F.T.) or the linear evaluation protocol (Lin.).
Linear Evaluation and Finetuning
In Table 2, we apply both the linear evaluation (Lin.) and finetuning (F.T.) protocols to our models across a diverse set of datasets consisting of CIFAR10 Krizhevsky (2009), CIFAR100, Flowers102Nilsback & Zisserman (2008), Food101Bossard et al. (2014), and iNaturalist 2021 (iNaturalist 2021 competition dataset, ). Our results show that the improvements achieved by sampling hard views we observed so far also transfer to other datasets.
Method Arch. VOC Det. COCO Segm.
SimSiam RN50 78.46 45.16
+ HVS RN50 79.06 46.51
Impr. +0.60 +1.35
DINO RN50 80.86 50.35
+ HVS RN50 80.51 50.39
Impr. -0.36 +0.04
Table 3: AP50 for object detection and instance segmentation with 100-ep. pretraining.
Object Detection and Instance Segmentation
For object detection, we use the VOC07+12 (Everingham et al., 2010) dataset using Faster R-CNN (Ren et al., 2015). We study instance segmentation on the COCO (Lin et al., 2014) dataset using Mask R-CNN (He et al., 2022). Table 3, where we report the AP50 performance, shows that the features learned with HVS also mostly transfer favourably to these tasks and outperform the SimSiam baseline. However, in case of DINO, we observe a slight performance gap on the detection task, which is the only instance in the entire paper where HVS did not outperform the baseline. This result warrants further investigation, particularly considering that this experiment was conducted with a single seed (due to limited availability of compute resources). More details and performance results on this task are provided in Appendix F.1.
5 Empirical Analysis of Hard View Selection
In this section, we discuss studies designed to shed light on the mechanisms underlying HVS. We address the following questions: 1) “Which can be observed that underlie the hard view selection?” 2) “Can a “manual” augmentation policy be inferred from these patterns?”, and 3) “What are the effects on empowering the adversary?”. For all experiments conducted here, we use our 100-epoch, ImageNet-pretrained SimSiam+HVS models with four sampled views.
5.1 Q1: Which Patterns Underlying the Hard View Selection Can be Observed?
When visually studying some of our examples and the views selected by HVS in Figures 2 and 6 (in the appendix), we notice that both geometric and appearance characteristics seem to be exploited. E.g., consider the brightness difference between the views of the first two rows in Fig. 2. Additionally, we observe a generally higher training loss (Fig. 7 in the appendix) indicative of an increased task difficulty. Another key observation is that our adversarial selection chooses view pairs that incorporate zooming in and out or an increased distance between the views (last row in Fig. 2).
Logging Augmentation Data
To assess these observations quantitatively, we logged relevant hyperparameter data during SimSiam training with HVS and conducted multiple studies which we will describe below. The data we logged include for each view the sampled parameters from the geometric and appearance data augmentation operations (such as the height/width and coordinates of the crops) as well as color distortion parameters (such as brightness and contrast, see Section D in the appendix for more details), as well as the loss and whether the view was selected or not. As evaluated metrics we chose the Intersection over Union (IoU), Relative Distance (image-wise normalized distance of the center points of view pairs), color distortion distance (the Euclidean distance between all four color distortion parameters), and the individual color distortion parameters.
Figure 3: Left: In over 40% of the cases, the adversarially selected view pair is also the one with the lowest IoU throughout SimSiam+HVS pretraining. We attribute the spike in the early phase to the random initialization of the embedding. Right: When using HVS (blue), a shift to smaller IoU values is visible when comparing to standard pretraining (orange). Both results are based on 3 seeds.
Importance of Augmentation Metrics
Given 300k such samples, we then used fANOVA (Hutter et al., 2014) to determine how predictive these metrics are. This resulted in the metric with the highest predictive capacity on the loss to be the IoU, explaining 15% of the variance in performance, followed by brightness with 5% (for more details see Fig. 9 in the appendix). The importance of IoU in HVS is further underpinned by the following observation: the fraction of view pairs selected by HVS, which are also the pairs with the lowest IoU among all six pairs (N=4), is over 40% (random: ${\sim}16.7\%$) during training. Moreover, when using HVS, a shift to smaller IoU values can be observed when comparing against standard pretraining. We visualize both findings in Fig. 3.
Taking a Closer Look at the Intersection over Union
We also examined the IoU value over the course of training in Fig. 4 (left). Here, we notice the average IoU value of selected view pairs increases slightly with training, possibly as a reaction of the pretrained model’s embedding representation to the hard view selection. Another pattern we observe is that the IoU value with HVS (Fig. 4 (left) in blue) is smaller and also less constant during training when compared against training without HVS (green). We believe this is due to the sample-wise and stateful nature of the adversarial selection as HVS chooses different IoU values from sample to sample and training state to training state. Lastly, we assessed the effect of the color augmentation on the pair selection. For this study, we sampled one set of color augmentations (as opposed to multiple, i.e. one each view) per iteration and applied it to all views. Only after identifying the hardest pair, we apply sampled data augmentations for each view individually. Consequently, the linear evaluation performance dropped by 0.3% on average. Interestingly, as we show in Fig. 4 (right), the fraction of selected pairs that are also the hardest pairs slightly increases in this case. One possible explanation for this is that it reflects the non-negligible role of color variation between views (as shown previously with the importance analysis), where HVS is given less leverage to increase hardness through a static appearance and instead, depends more on leveraging the IoU.
5.2 Q2: Can a Manual Augmentation Policy be Inferred?
IoU Policy Type SimSiam DINO
Baseline (B) 68.20 73.50
B+range(0.3-0.35) -0.80 -1.47
B+range(0.3-0.35)+alternating +0.10 -0.45
B+range(0.4-0.6) +0.55 -0.40
B+range(0.4-0.6)+alternating +0.25 -0.20
B+range(0.1-0.8) -33.95 -1.50
B+range(1.0-0.1) +0.07 -
Table 4: Top-1 lin. eval. accuracies for the manual IoU policy (averaged over 2 seeds).
Since harder pretraining tasks seem beneficial, and based on the observations made in Q1, a natural question arises: is it possible to mimic the adversarial selection with a manually scripted augmentation policy? Such a policy, if effective, would replace HVS and lower the computational cost by avoiding the additional forward passes. Since the IoU seems to play an important role, below, we study several possible ways for constructing a simple manual augmentation policy based on IoU.
Deriving an Augmentation Policy
We implemented the following algorithm based on rejection sampling in the augmentation pipeline: we linearly approximate the IoU value curve from Fig. 4 (left; in blue) with a start value of 0.30 and end value 0.35 (ignoring the dip in the early phase) to receive an upper bound for the IoU in each training iteration. For each iteration, we then check if the pairs exceed the IoU upper bound value and if so, we reject the pair and re-sample a new pair (max. 512 times). This ensures that only view pairs are sampled that employ at least a certain task difficulty (by means of a small enough IoU). We varied different hyperparameters (e.g., IoU start/end ranges; inverse schedules; alternating between the IoU schedule and the standard augmentation every other iteration; as well as warmup, i.e. no IoU policy for the first 10 epochs). Training both SimSiam (ResNet-50) and DINO (ViT-S/16; applied to all global and local heads) models for 100 epochs yielded performance decreases or insignificant improvements, as can be seen in Table 4. Moreover, we see an increased risk of model collapse when setting start values too small which can be interpreted as a too hard task. These results indicate that developing a manual policy based on metrics in pixel space such as the IoU is non-trivial. In contrast, we believe that HVS is effective since it 1) operates on a similarity level of latent embeddings that may be decorrelated from the pixel space and 2) has access to the current model state.
Figure 4: Left: The average IoU of view pairs selected by SimSiam+HVS (blue) compared against the default SimSiam training (green). Right: Using static color augmentation for all pairs before the selection increases the dependency on the IoU.
Assessing the Difficulty of Predicting the Pair Selected by HVS
To further validate the previous result and to ascertain an upper limit on the performance for predicting the hardest pair, we conducted a second experiment. We fitted a gradient boosting classification tree (Friedman, 2001) to predict the selected view pair conditioned on all SimSiam hyperparameter log data from Q1 (see above) except for the flag that indicates whether a view was selected. As training and test data, we used the logs from two seeds (each 300k samples) and the logs from a third seed, respectively. We also tuned hyperparameters on train/valid splits and applied a 5-fold CV. However, the resulting average test performance in all scenarios never exceeded 40%, indicating that it is indeed challenging to predict the hardness of views based on parameter-level data.
Consequently, we believe these results support the presumption that deriving a policy for controlling and increasing hardness based on geometric and appearance parameters is non-trivial. Furthermore, our analysis suggests that such a policy must function on a per-sample basis and have access to the current model state which is the case when applying HVS.
5.3 Q3: What are the Effects of Empowering the Adversary?
It is well known that adversarial learning can suffer from algorithmic instability (Xing et al., 2021), e.g. by giving an adversary too much capacity. We now further explore the space of adversarial capacity for selecting hard views to gain a better understanding of its impact on contrastive learning.
Optimizing an Adversarial Learner for the View Generation
In this experiment, we briefly explore adversarially learning a network to output the transformation matrix for view generation. We optimize a Spatial Transformer Network (STN) (Jaderberg et al., 2015) jointly with the DINO objective and a ViT-tiny. Using the inverse of the DINO loss, we train the STN to output the geometric transformation matrices (each 6 parameters) for generating views given an input image that are used for standard CL training with DINO. We employ and test various transformation types (e.g., rotation, translation, etc.) that influence the capacity of the adversary. We discuss this experiment in greater detail in Appendix E and briefly summarize our findings here: applying an adversary in this way is very challenging due to the sensitivity of CL on the data augmentation and the dynamics of adversarial learning, often leading to either minor improvements or model collapses as we have observed when employing a too strong manual IoU policy in Q1.
Increasing the Number of Views
In our initial experiments, we explored variations in the number of sampled views $N$ with SimSiam and HVS. As can be seen in Fig. 8 in the appendix, while $N=8$ views still outperform the baseline in terms of linear evaluation accuracy, it is slightly worse than using $N=4$ views (-0.05% for 100 epochs and -0.14% for 200 epochs pretraining on linear evaluation). We interpret this result as the existence of a “sweet spot” in setting the number of views, where, in the limit, a higher number of views corresponds to approximating the adversarial learner described above. This would allow HVS to select progressively harder views that quickly results in very hard tasks and consequently in model collapse and performance deterioration.  [^1] For completeness, we also experimented with a cooperative HVS (see Appendix F.2 for details).
Figure 5: Comparing DINO with DINO+HVS on CIFAR100 for longer trainings.
Longer Training
Despite the inherent overhead associated with additional forward passes of HVS, our results clearly show that using HVS is also beneficial for long trainings (see 300-epoch improvements in Table 1). To delve deeper into the behavior of longer trainings under our computational constraints, we conduct several training runs with progressively higher epochs on the CIFAR100 dataset, using the same initial learning rate and a cosine schedule. The comparative analysis in Fig. 5 shows, except for some overfitting, no evidence of diminishing returns as the improvement remains consistent throughout training nor signs of instability (e.g., collapse) due to hard tasks.
6 Conclusion
In this study, we presented HVS, a new data augmentation and learning strategy for contrastive learning designed to challenge pretrained models with more demanding samples. This method, while straightforward in its design, has proven to be a powerful tool, pushing the boundaries of the traditional random view generation in contrastive learning. When combined with methods like DINO, SimSiam, and SimCLR, HVS consistently showcased improvements in linear evaluation and a diverse set of transfer tasks. As the landscape of self-supervised learning continues to evolve, methods like HVS underscore the potential for further advancements in the field. By continuously challenging models with harder samples and refining augmentation techniques, we anticipate even greater strides in model performance and generalization in future research endeavors. While HVS incurs a computational overhead due to its additional forward passes, our preliminary experiments indicate no signs of diminishing returns with extended training durations. This suggests that HVS is likely to continue to outperform baselines regardless of the training length. Looking ahead, we believe there is ample room for refining and expanding HVS, such as further investigating the number of sampled views based on the current learning state, enhancing efficiency through checkpointing activations, bypassing forwarding of similar pairs, or augmentation distributions that allow controlling task difficulty.
Appendix A Examples Sampled by HVS
Figure 6: We depict row-wise ten example images from the ImageNet train set along with their sampled views with a finished, 100-epoch trained SimSiam ResNet50 model. Left: original image with the overlaid randomly sampled crops (colored dashed rectangles). Right: All views after applying resizing and appearance augmentations. The pair that is selected adversarially by HVS is highlighted in solid lines, eg. View 1 and View 4 in the first row.
Appendix B Training Loss
Figure 7: The training loss over 100 epochs. Comparing the DINO vanilla method with DINO + HVS. The spike and drop in the loss curve of DINO is caused by freezing the last layer in the first epoch which was proposed by the authors as a strategy to enhance downstream performance. For HVS we can only see a drop and no spike. We believe this is because HVS exposes the model to hard views from the beginning of training (i.e. the loss is immediately maximized).
Appendix C Effect of More Views on Linear Evaluation Performance
Figure 8: Setting the number of views too high can result in performance deterioration. This shows that diminishing returns exist, likely because the adversary becomes too strong, resulting in a too hard learning task.
Appendix D Assessing the Importance of Metrics with fANOVA
To assess the importance of various metrics on the training loss, we apply fANOVA Hutter et al. (2014) on data that we logged during training with HVS. We used 300k samples that contain the following sampled parameters from the geometric and appearance data augmentation operations for each view: all random resized crop parameters (height and width of the original image, coordinates of crop corners and height and width of the crop), all Colorjitter (color distortion) strengths (brightness, contrast, saturation, hue), grayscale on/off, Gaussian blurring on/off, horizontal flip on/off, loss, and if the crop was selected or not. The metrics we chose are Intersection over Union (IoU), Relative Distance (sample-wise normalized distance of the center points of crop pairs), color distortion distance (the Euclidean distance between all four color distortion operation parameters, i.e. brightness, contrast, saturation, hue), and the individual color distortion parameters of the Colorjitter operation. As can be seen in Fig. 9, the metric with the highest predictive capacity on the loss is the IoU with an importance of 15.2% followed by brightness with 5.1%. The relative distance has an importance of 3.3%, the colorjitter distance 2.3%, the contrast 1.6%, the saturation 1.4%, hue 0.6%, and all parameters jointly 1.7%.
Figure 9: Application of fANOVA Hutter et al. (2014) on logged training data to determine metrics with high predictive capacity on the train loss.
Appendix E Adversarial Learner for the View Generation
We use Spatial Transformer Network (STN) (Jaderberg et al., 2015) to allow generating views by producing 6D transformation matrices (allowing translating, rotating, shearing, scaling, affine transformations and combinations thereof) in a differentiable way since most common augmentations are not off-the-shelf differentiable. As described, we train it alongside the actual pretrained network using the same (inverted) objective. For our experiments we use DINO with multi-crop, i.e. 2 global and 8 local heads. As STN we use a small CNN with a linear layer for outputting the 10*6D transformation matrices. In this scenario, we use a ViT-tiny/16 with a 300 epoch pretraining on CIFAR10 with a batch size of 256. All other hyperparameters are identical to the ones reported in the DINO paper.
Figure 10 visualizes the procedure. The STN takes the raw image input and generates a number of transformation matrices that are applied to transform the image input into views. These views are then passed to the DINO training pipeline. Both networks are trained jointly with the same loss function. DINO is trained with its original contrastive objective, where the STN is trained by inverting the gradient after the DINO during backpropagation.
{tikzpicture}
[…]
Figure 10: Illustration of adversarial learning with a Spatial Transformer Network (STN) jointly with contrastive learning (here: DINO).
As mentioned previously, the STN, without using auxiliary losses, starts zooming in and generating single-color views. To counteract this behavior, we experimented with different penalties on the transformation matrices produced by the STN. For instance, in order to limit the zooming pattern, we can use the determinants of the sub-matrices of the transformation matrix to penalize based on the area calculated and apply a regression loss (e.g. MSE). We refer to this type of penalty as Theta Crops Penalty (TCP). Additionally, we also restrict its parameters to stay within a sphere with different parameters for local and global crops. Next to determinant-based penalty losses, we also experimented with other penalty functions such as the weighted MSE between the identity and the current transformation matrix or penalties based on histograms of the input image and generated views after applying the transformation. To avoid strong uni-dimensional scaling behavior, we also implemented restricting scaling in a symmetric way (i.e. applied to both x and y dimensions) and refer to this as scale-sym.. We report our best results in Table 5 which are all TCP-based. As can be seen, no setting is able to outperform the baseline. Our best score was achieved with translation-scale-symmetric which is very similar to random cropping. When removing the symmetries in scaling, the performance drops further. Removing a constraint adds one transformation parameter and therefore one dimension. This can be seen as giving more capacity to the adversarial learner which in turn can make the task significantly harder. Similarly, when adding rotation, the performance drops further and in part drastically. This is on the one hand due to the penalties not being fully able to restrict the output of the STN. On the other, the task of extracting useful information from two differently rotated crops is even harder, and learning spatial invariance becomes to too challenging. All in all, we experienced two modes: either the STN is too restricted, leading to static output (i.e. independent of image content, the STN would produce constant transformation matrices) or the STN has too much freedom, resulting in extremely difficult tasks. See Fig. 11 for an example on the former behavior.
Figure 11: Example for static output behavior of the STN.
Mode Penalty Lin. F.T.
baseline - 86.1 92.7
translation-scale-sym. TCP 83.7 90.3
translation-scale TCP 82.8 89.7
rotation-translation TCP 56.7 -
rotation-translation-scale TCP 31.7 -
rotation-translation-scale-sym. TCP 77.6 -
affine TCP 78.3 83.5
Table 5: Linear evaluation and finetuning classification performance on CIFAR10. Top-1 accuracy on the validation set of CIFAR10 for our best results reported with different STN transformation modes.
Appendix F Additional Results
F.1 Object Detection and Instance Segmentation
For our object detection and instance segmentation analysis experiment, we report additional results in Table F.1 based on the AP metric. We use the C4 backbone variant (Wu et al., 2019) and finetuning with the 1x schedule.
Method Arch. VOC07+12 Object Det. COCO Instance Segm.
$\mathbf{AP_{all}}$ $\mathbf{AP_{50}}$ $\mathbf{AP_{75}}$ $\mathbf{AP_{mask}}$ $\mathbf{AP_{mask50}}$ $\mathbf{AP_{mask75}}$
SimSiam RN50 52.05 78.46 56.88 27.16 45.16 28.48
+ HVS RN50 52.91 79.06 57.55 27.93 46.51 29.36
Improvement +0.86 +0.60 +0.67 +0.77 +1.35 +0.88
DINO RN50 53.66 80.86 58.60 30.25 50.35 31.78
+ HVS RN50 52.86 80.51 58.21 30.02 50.39 31.52
Improvement +0.80 -0.36 -0.39 -0.23 +0.04 -0.27
Table 6:
F.2 Easy View Selection
To investigate the effect of a cooperative, i.e. easy pair selection, we conducted a small experiment. Instead of selecting the pair yielding the worst loss, we inverted the objective and selected the pair with the best loss. As expected, this led to model collapses with a linear eval. performance of 0.1%. This result is in line with previous findings that highlight the importance of strong augmentations in CL.
Appendix G Hyperparameters
G.1 Evaluations on ImageNet
G.1.1 DINO
For DINO, we report the ViT pretraining hyperparameters in Table 7. For ResNet-50, we use the same hyperparameters. Note, for HVS we limit the total number of comparisons to 128 across all heads. Linear evaluation is executed for 100 epochs and we use a learning rate of $0.001$, SGD optimizer (AdamW (Loshchilov & Hutter, 2019) during pretraining), a batch size of 1024, a momentum of $0.9$, and no weight decay.
Hyperparameter Value Hyperparameter Value
architecture vit_small epochs: 100
img_size 224 warmup_epochs: 10
patch_size 16 freeze_last_layer: 1
out_dim 65536 lr: 0.0005
norm_last_layer true min_lr: 1.0e-06
momentum_teacher 0.996 optimizer: AdamW
use_bn_in_head false weight_decay: 0.04
teacher_temp 0.04 weight_decay_end: 0.4
warmup_teacher_temp 0.04 global_crops_scale: 0.4, 1.0
warmup_teacher_temp_epochs 0 global_crops_size: 224
fp16 true local_crops_number: 8
batch_size 512 local_crops_scale 0.05, 0.4
clip_grad 3.0 local_crops_size: 96
drop_path_rate 0.1
Table 7: Pretraining ImageNet hyperparameters for the runs with DINO. For 300 epochs, we use a batch size of 1024.
G.1.2 SimSiam
In Table 8, we report the ResNet-50 pretraining hyperparameters. Linear evaluation is executed for 90 epochs (as reported by the SimSiam authors) and we use a learning rate of 0.1, LARS optimizer (You et al., 2017), a batch size of 4096, and no weight decay.
Hyperparameter Value
architecture resnet50
batch_size 512
blur_prob 0.5
crops_scale 0.2, 1.0
crop_size 224
feature_dimension 2048
epochs 100
fix_pred_lr true
lr 0.05
momentum 0.9
predictor_dimension 512
weight_decay 0.0001
optimizer SGD
Table 8: Pretraining ImageNet hyperparameters for the runs with SimSiam. For 300 epochs, we use a batch size of 1024.
G.1.3 SimCLR
We report the ResNet-50 pretraining hyperparameters for SimCLR in Table 9. Linear evaluation is executed for 90 epochs with a learning rate 0.1, SGD optimizer, batch size of 4096 and no weight decay.
Hyperparameter Value
architecture resnet50
proj_hidden_dim 2048
out_dim 128
use_bn_in_head true
batch_size 4096
optim LARS
lr 0.3
sqrt_lr false
momentum 0.9
weight_decay 1e-4
epochs 100
warmup_epochs 10
zero_init_residual true
Table 9:
G.2 Transfer to Other Datasets and Tasks
For linear evaluation on the transfer datasets, we simply used the same hyperparameters for linear evaluation on ImageNet (DINO and SimSiam respectively). For finetuning DINO ViT-S/16, we used the hyperparameters reported in Table 10 and for SimSiam ResNet-50 we used the hyperparameters in Table 11
Hyperparameter CIFAR10 CIFAR100 Flowers102 iNat 21 Food101
lr 7.5e-6 7.5e-6 5e-5 5e-5 5e-5
weight_decay 0.05 0.05 0.05 0.05 0.05
optimizer AdamW AdamW AdamW AdamW AdamW
epochs 300 300 300 100 100
batch_size 512 512 512 512 512
Table 10: Finetuning hyperparameters for DINO ViT-S/16.
Hyperparameter CIFAR10 CIFAR100 Flowers102 iNat 21 Food101
lr 7.5e-6 5e-6 5e-4 7e-5 5e-6
weight_decay 0.05 0.05 0.05 0.05 0.05
optimizer AdamW AdamW AdamW AdamW AdamW
epochs 300 300 300 100 100
batch_size 512 512 512 512 512
Table 11: Finetuning hyperparameters for SimSiam and ResNet-50.
G.3 Object Detection and Instance Segmentation
We have used the Detectron2 library (Wu et al., 2019) for object detection and instance segmentation. We followed the public codebase from MoCo (He et al., 2020) (like SimSiam) for all entries. Due to limited compute resources we changed the batch size. All parameters that differ from MoCo are reported in the Table 12. The pretrained models are finetuned end-to-end on the target datasets. All methods are based on 100-epoch pre-training on ImageNet.
VOC07+12 Object Det. Coce Inst. Segm.
Hyperparameter Value Value
batch_size 8 8
lr 0.01 0.01
steps 48000 90000
Table 12: Hyperparameters for VOC object detection and COCO instance segmentation.
Appendix H Hard View Selection Objectives
H.1 SimCLR
In this section, we are going to introduce the application of HVS with the SimCLR objective. Assume a given set of images $\mathcal{D}$, an image augmentation distribution $\mathcal{T}$, a minibatch of $M$ images $\mathbf{x}=\left\{x_{i}\right\}_{i=1}^{M}$ sampled uniformly from $\mathcal{D}$, and two sets of randomly sampled image augmentations $A=\left\{t_{i}\sim\mathcal{T}\right\}_{i=1}^{M}$ and $B$ sampled from $\mathcal{T}$. We apply $A$ and $B$ to each image in $\mathbf{x}$ resulting in $\mathbf{x}^{A}$ and $\mathbf{x}^{B}$. Both augmented sets of views are subsequently projected into an embedding space with $\mathbf{z}^{A}=g_{\theta}(f_{\theta}(\mathbf{x}^{A}))$ and $\mathbf{z}^{B}=g_{\theta}(f_{\theta}(\mathbf{x}^{B}))$ where $f_{\theta}$ represents an encoder (or backbone) and $g_{\theta}$ a projector network. Contrastive learning algorithms then minimize the following objective function:
$\mathcal{L(\mathbf{\mathcal{T},x;\theta})}=-\log\frac{\exp(\text{sim}(\mathbf{z}_{i}^{A},\mathbf{z}_{i}^{B})/\tau)}{\sum_{i\neq j}\exp(\text{sim}(\mathbf{z}_{i}^{A},\mathbf{z}_{j}^{B})/\tau)}$ (3)
where $\tau$ denotes a temperature parameter and sim a similarity function that is often chosen as cosine similarity. Intuitively, when optimizing $\theta$, embeddings of two augmented views of the same image are attracted to each other while embeddings of different images are pushed further away from each other.
To further enhance the training process, we introduce a modification to the loss function where instead of having two sets of augmentations $A$ and $B$, we now have ”N” sets of augmentations, denoted as $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{N}\}$. Each set $A_{i}$ is sampled from the image augmentation distribution $\mathcal{T}$, and applied to each image in $\mathbf{x}$, resulting in ”N” augmented sets of views $\mathbf{x}^{A_{1}},\mathbf{x}^{A_{2}},\ldots,\mathbf{x}^{A_{N}}$.
Similarly, we obtain $N$ sets of embeddings $\mathbf{z}^{A_{1}},\mathbf{z}^{A_{2}},\ldots,\mathbf{z}^{A_{N}}$ through the encoder and projector networks defined as:
$\mathbf{z}^{A_{i}}=g_{\theta}(f_{\theta}(\mathbf{x}^{A_{i}})),\quad i=1,2,\ldots,N$
We then define a new objective function that seeks to find the pair of augmented images that yield the highest loss. The modified loss function is defined as:
$\mathcal{L}_{\text{max}}(\mathcal{T},\mathbf{x};\theta)=\max_{k,l:k\neq l}\mathcal{L}(\mathcal{T},\mathbf{x};\theta)_{kl}$
where
$\mathcal{L}(\mathcal{T},\mathbf{x};\theta)_{kl}=-\log\frac{\exp(\text{sim}(\mathbf{z}_{k}^{A_{k}},\mathbf{z}_{k}^{A_{l}})/\tau)}{\sum_{i\neq j}\exp(\text{sim}(\mathbf{z}_{i}^{A_{k}},\mathbf{z}_{j}^{A_{l}})/\tau)}$
and $k,l\in\{1,2,\ldots,N\}$ and $i,j\in\{1,2,\ldots,M\}$.
In this formulation, we iterate over all possible pairs of augmented images derived from the same original image across all ”N” augmentations, and compute the loss for each pair. The pair that gives the highest loss is then used in the optimization process. It is important to note that the temperature parameter $\tau$ and the similarity function remain as defined in the original loss function.

### v4

@@
-Hard View Selection for Contrastive Learning
+Hard View Selection for Self-Supervised Learning
 Abstract
-Many Contrastive Learning (CL) methods train their models to be invariant to different “views” of an image input for which a good data augmentation pipeline is crucial.
-While considerable efforts were directed towards improving pre-text tasks, architectures, or robustness (e.g., Siamese networks or teacher-softmax centering), the majority of these methods remain strongly reliant on the random sampling of operations within the image augmentation pipeline, such as the random resized crop or color distortion operation.
-In this paper, we argue that the role of the view generation and its effect on performance has so far received insufficient attention.
-To address this, we propose an easy, learning-free, yet powerful Hard View Selection (HVS) strategy designed to extend the random view generation to expose the pretrained model to harder samples during CL training.
-It encompasses the following iterative steps: 1) randomly sample multiple views and create pairs of two views, 2) run forward passes for each view pair on the currently trained model, 3) adversarially select the pair yielding the worst loss, and 4) run the backward pass with the selected pair.
-In our empirical analysis we show that under the hood, HVS increases task difficulty by controlling the Intersection over Union of views during pretraining.
-With only 300-epoch pretraining, HVS is able to closely rival the 800-epoch DINO baseline which remains very favorable even when factoring in the slowdown induced by the additional forwards of HVS.
-Additionally, HVS consistently achieves accuracy improvements on ImageNet between 0.55% and 1.9% on linear evaluation and similar improvements on transfer tasks across multiple CL methods, such as DINO, SimSiam, and SimCLR.
+Many Self-Supervised Learning (SSL) methods train their models to be invariant to different “views” of an image and considerable efforts were directed towards improving pre-text tasks, architectures, or robustness.
+However, most SSL methods remain reliant on the random sampling of operations within the image augmentation pipeline, such as the random resized crop operation.
+We argue that the role of the view generation and its effect on performance has so far received insufficient attention.
+To address this, we propose an easy, learning-free, yet powerful Hard View Selection (HVS) strategy designed to extend the random view generation to expose the pretrained model to harder samples during SSL training.
+It encompasses the following iterative steps: 1) randomly sample multiple views and create pairs of two views, 2) run forward passes for each view pair on the currently trained model, 3) adversarially select the pair yielding the worst loss depending on the current model state, and 4) run the backward pass with the selected pair.
+As a result, HVS consistently achieves accuracy improvements between 0.91% and 1.93% on ImageNet linear evaluation and similar improvements on transfer tasks across DINO, SimSiam, iBOT and SimCLR.
+We provide studies to shed light on the inner workings and show that, by naively using smaller resolution images for the selection step, we can significantly reduce the computational overhead while retaining performance.
+Surprisingly, even when accounting for the computational overhead incurred by HVS, we achieve performance gains between 0.52% and 1.02% and closely rival the 800-epoch DINO pretraining with only 300 epochs.
 1 Introduction
@@
 In this paper, we seek to extend the random view generation to expose the pretrained model to harder samples during CL training.
-Figure 2: Example images (left) and sampled views (right) after resizing and appearance transformations. Pairs selected adversarially by HVS are shown with a solid frame.
-Supervised Learning works that address this shortcoming in a distant manner and which attempt to leverage task difficulty include Spatial Transformer Networks (Jaderberg et al., 2015)and Adversarial AutoAugment (Zhang et al., 2020), which either directly learn geometric transformations or generate augmentation policies that yield hard samples for more robust feature learning.
-However, these approaches have not been used successfully for unsupervised learning and Adversarial AutoAugment is limited in leveraging the geometric perspective for view hardness control as it does not include RRC in its search space.
-Furthermore, and in contrast to what we propose, integrating any of these methods into existing contrastive learning pipelines is non-trivial.
-Contrastive learning approaches that are related to our work and aim at controlling view hardness include (Shi et al., 2022) and (Tamkin et al., 2021) which propose masking or perturbing input images through adversarial models.
-Another approach is MADAug (Hou et al., 2023) which use a cooperative setting to learn policies yielding progressively harder samples based on a validation loss which is usually unavailable in contrastive learning.
-A work that takes a similar perspective like us in leveraging the information content of views for contrastive learning is (Tian et al., 2020b).
-The authors unify the data augmentation of several CL methods under the lens of mutual information and learn view generators under this premise.
-While these works offer important insights into view optimality, they usually require learning additional models or non-trivial changes to the pipeline and thus, are not easily integrated into existing CL methods.
-In contrast, we propose a fully learning-free, easy-to-integrate approach that applies the following simple strategy: we sample a few views and forward each pair through the currently trained model.
+Figure 2: Examples (left) and sampled views (right) after resizing and appearance transformations. Pairs selected by HVS are shown with a solid frame.
+Supervised Learning works that address this shortcoming and which attempt to leverage task difficulty include Spatial Transformer Networks (Jaderberg et al., 2015) and Adversarial AutoAugment (Zhang et al., 2020), which either directly learn geometric transformations or generate augmentation policies that yield hard samples.
+However, these approaches have not been used successfully for unsupervised learning and Adversarial AutoAugment does not include RRC in its search space.
+In contrast to what we propose, integrating any of these methods into existing self-supervised learning pipelines is non-trivial.
+Related SSL approaches exist that aim at controlling view hardness through adversarial Shi et al.
+(2022); Tamkin et al.
+(2021) or cooperative Hou et al.
+(2023) models.
+A work that takes a similar perspective like us in leveraging the information content of views for SSL is (Tian et al., 2020b).
+They discover sweet spots for data augmentation hyperparameters that are widely used and adversarially learn flow-based models that generate novel color space-views which are then used for contrastive learning.
+While these works offer important insights into view optimality, they usually require learning additional models prior to the actual SSL training or non-trivial changes to the pipeline and thus, are not easily integrated into existing SSL methods.
+In contrast, we propose a fully learning-free and easy-to-integrate approach.
+Importantly, and different to Tian et al.
+(2020b), our approach leverages the current model state and operates on a per-sample basis during pretraining.
+It applies the following simple strategy: we sample a few views and forward each pair through the currently trained model.
 Next, we identify the pair that yields the worst loss among all pairs, use it for the backward pass, and repeat (see Fig.
 1).
-While being agnostic to any contrastive loss, this strategy exposes the pretrained model to samples that the model at any given point in the training trajectory finds challenging.
+While being agnostic to any contrastive and non-contrastive loss, this strategy exposes the pretrained model to samples that the model at any given point in the training trajectory finds challenging.
 As our results show, sampling only four views are sufficient to achieve noticeable improvements.
 Overall, our contributions can be summarized as follows:
-  • We propose Hard View Selection (HVS), a new learning method complementing contrastive learning that is easy to integrate and extends the broadly-used random view generation to automatically expose the model to harder samples during pretraining.
+  • We propose Hard View Selection (HVS), a new learning method complementing SSL that is easy to integrate and extends the broadly-used random view generation to automatically expose the model to harder samples during pretraining.
 The only requirement of HVS is that it needs to be able to compute sample-wise losses;
-  • We demonstrate the effectiveness and compatibility of our approach using ImageNet (Deng et al., 2009) pretraining across popular and widely-adopted contrastive learning methods that cover a diverse range of objectives such as SimSiam (Chen & He, 2021), DINO (Caron et al., 2021), SimCLR (Chen et al., 2020a);
-  • HVS improves all baselines noticeably compared to using the common view generation and we show this on a diverse set of transfer datasets and tasks, including linear evaluation, finetuning, object detection and segmentation;
-  • We present insights into the underlying mechanisms of HVS.
-One key observation is that harder samples are achieved by scheduling the Intersection over Union conditioned on the current model state and image input.
-We also provide further results on the effects of changing the number of sampled views, color augmentation and studies on longer training.
+  • We demonstrate the effectiveness and compatibility of our approach using ImageNet (Deng et al., 2009) pretraining across popular and widely-adopted SSL methods that cover a diverse range of objectives such as SimSiam (Chen & He, 2021), DINO (Caron et al., 2021), iBOT (Zhou et al., 2021), and SimCLR (Chen et al., 2020a);
+  • HVS improves all baselines noticeably compared to using the common view generation, and we show this on a diverse set of transfer datasets and tasks, including linear evaluation, finetuning, object detection and segmentation;
+  • We present insights into the underlying mechanisms of HVS such as the effect of changing the number of sampled views or color augmentation.
 We make our PyTorch (Paszke et al., 2019) code, models, and all used hyperparameters publicly available under https://anonymous.4open.science/r/hard-view-selection/.
 2 Related Work
-2.1 Contrastive Learning in SSL
+2.1 Contrastive and Non-Contrastive Learning in SSL
 The core idea behind the contrastive learning framework is to learn image representations by contrasting positive pairs (two views of the same image) against negative pairs (two views of different images) (Hadsell et al., 2006).
 To work well in practice and to prevent model collapse, contrastive learning methods often require a large number of negative samples (Wu et al., 2018; van den Oord et al., 2018; Chen et al., 2020a; He et al., 2020; Tian et al., 2020a; Chen et al., 2020b) stored in memory banks (Wu et al., 2018; He et al., 2020) or, for instance, in the case of SimCLR, implicitly in large batches (Chen et al., 2020a).
-Non-contrastive approaches, such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), DINO (Caron et al., 2021) and others (Zbontar et al., 2021; Caron et al., 2020; Ermolov et al., 2021), are able to only use positive pairs without causing model collapse but rely on other techniques, such as Siamese architectures, whitening of embeddings, clustering, maximizing the entropy of the embeddings, momentum encoders and more.
+Non-contrastive approaches, such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), DINO (Caron et al., 2021) and others (Zbontar et al., 2021; Caron et al., 2020; Ermolov et al., 2021), are able to only use positive pairs without causing model collapse but rely on other techniques, such as Siamese architectures, whitening of embeddings, clustering, maximizing the entropy of the embeddings, momentum encoders, and more.
 2.2 Optimizing for Hard Augmentations and Views in SSL
@@
 Most related to our work among these approaches is Adversarial AutoAugment (AAA) (Zhang et al., 2020) that proposes to generate augmentation policies yielding hard samples by learning with an adversarial supervised objective.
-In a similar, but cooperative setting, MADAug (Hou et al., 2023) learns policies yielding progressively harder samples based on a validation loss.
+In a similar, but cooperative setting, MADAug (Hou et al., 2023) learns policies yielding progressively harder samples.
 In contrast to our work, however, AAA and MADAug do not include the random resize crop operation into their search spaces and thus, are limited in leveraging the geometric perspective to control view hardness.
@@
 (2020b) leverage, like us, the view content for better performance.
-Under the lens of mutual information (MI), they unify several data augmentation CL methods and derive an MI-based objective to learn view generators for unsupervised and semi-supervised learning.
+Under the lens of mutual information (MI), they derive an objective to adversarially learn flow-based view generators for contrastive learning.
 While these methods can complement our approach, they are less easily integrated into existing contrastive pipelines due to the need for adversarial network training.
+Moreover, Tian et al.
+(2020b) does not use the current model state to control learning hardness and, while their studies incorporate distance metrics, they do not consider the Intersection over Union (IoU) metric which we extensively make use of.
 3 Method
@@
 Many different contrastive learning (He et al., 2020) objectives exist, each characterized by variations stemming from design choices, such as by the use of positive and negative samples or asymmetry in the encoder-projector network structure.
-For simplicity of exposure, we will introduce our approach based on the SimSiam objective Chen & He (2021), but we do note that our method can be used with any other contrastive learning objective that allows the computation of a sample-wise losses.
+For simplicity of exposure, we will introduce our approach based on the SimSiam objective Chen & He (2021), but we do note that our method can be used with any other discriminative SSL objective that allows the computation of a sample-wise losses.
 SimSiam works as follows.
@@
 We then select all pairs that yielded the highest loss to form the new hard minibatch of augmented sets $\textbf{x}^{A_{k_{*}}}$ and $\textbf{x}^{B_{l_{*}}}$, discard the other pairs and use the hard minibatch for optimization.
-As shown in Algorithm 12, we repeat this process in each training iteration.
-This modification introduces a more challenging learning scenario in which the model is encouraged to learn more discriminative features by being exposed to harder views.
-While the additional forward passes come with a certain overhead (about a factor of 1.55$\times$ for SimSiam), as we will demonstrate in our experiments, the focus on harder views leads to consistent improvements.
 1:Input: Number of views $N\geq 2$, batch size $M$,
@@
 5:     Create augmented views: $\textbf{x}_{i}^{A}=\left\{t_{n}(x_{i})\right\}_{n=0}^{N}$
-6:     Create all $\binom{N}{2}$ possible pairs $\textbf{x}_{i}^{A_{k}}\times\textbf{x}_{i}^{A_{l}}$, $k\neq l$
-7:     For all pairs $(x_{i}^{A_{k}},x_{i}^{A_{l}})$ forward through $f_{\theta}$ and $g_{\theta}$
+6:     Forward all views through $f_{\theta}$ and $g_{\theta}$ to get $z$ and $h$
+7:     Create all $\binom{N}{2}$ possible pairs $\textbf{x}_{i}^{A_{k}}\times\textbf{x}_{i}^{A_{l}}$, $k\neq l$
 8:     Select the hard pair $(x_{i}^{A_{k_{*}}},x_{i}^{A_{l_{*}}})$ maximizing Eq. 2
@@
 Algorithm 1 Contrastive Learning with HVS
+As shown in Algorithm 12, we repeat this process in each training iteration.
+This modification introduces a more challenging learning scenario in which the model is encouraged to learn more discriminative features by being exposed to harder views.
+While the additional forward passes come with a certain overhead (about a factor of 1.55$\times$ for SimSiam), as we will demonstrate in our experiments, the focus on harder views leads to consistent improvements.
 While we exemplified the integration of HVS with the SimSiam objective, integrating it into other contrastive methods is as straightforward.
 The only requirement of HVS is to be able to compute sample-wise losses (in order to select the views with the highest loss).
-In our experiments section and in addition to SimSiam, we study the application of HVS to the objectives of DINO and SimCLR (see also Appendix H.1 for a formal exemplary definition for the integration of HVS into SimCLR).
+In our experiments section and in addition to SimSiam, we study the application of HVS to the objectives of DINO and SimCLR (see also Appendix I.1 for a formal exemplary definition for the integration of HVS into SimCLR).
 3.3 Implementation and Evaluation Protocols
@@
 DINO RN50 71.93 66.28 75.25 69.53
-+ HVS RN50 72.87 67.33 75.80 70.11
-Improvement +0.94 +1.05 +0.55 +0.58
++ HVS RN50 72.87 67.33 75.65 70.05
+Improvement +0.94 +1.05 +0.40 +0.52
 SimSiam RN50 68.20 57.47 70.35 61.40
@@
 Improvement +1.93 +1.85 +1.70 +1.66
+iBOT ViT-S/16 69.50 62.82 72.48 66.60
++ HVS ViT-S/16 70.41 63.05 73.99 67.12
+Improvement +0.91 +0.23 +1.51 +0.52
 Table 1: Average top-1 linear and k-NN classification accuracy on the ImageNet validation set for 100 and 300-epoch pretrainings. For SimCLR we run 200 instead of 300 epochs.
@@
 Here, we discuss our main results on image classification, object detection, and segmentation tasks.
-All results are self-reproduced using the baseline code and available hyperparameters, and are averaged over multiple seeds.
+All results are self-reproduced using the baseline code and available hyperparameters, and are averaged over 3 seeds (iBOT over 2 seeds).
 All hyperparameters are reported in Appendix G.
@@
 Impr. -0.36 +0.04
-Table 3: AP50 for object detection and instance segmentation with 100-ep. pretraining.
+Table 3: AP50 for object detection and instance segmentation.
 Object Detection and Instance Segmentation
@@
 We study instance segmentation on the COCO (Lin et al., 2014) dataset using Mask R-CNN (He et al., 2022).
-Table 3, where we report the AP50 performance, shows that the features learned with HVS also mostly transfer favourably to these tasks and outperform the SimSiam baseline.
+Table 3, where we report the AP50 performance, shows that the features learned with HVS also mostly transfer favourably to these tasks and outperform the SimSiam baseline with a 100-ep.
+pretraining.
 However, in case of DINO, we observe a slight performance gap on the detection task, which is the only instance in the entire paper where HVS did not outperform the baseline.
-This result warrants further investigation, particularly considering that this experiment was conducted with a single seed (due to limited availability of compute resources).
+This result warrants further investigation, as it was conducted with a single seed (due to limited compute resources).
 More details and performance results on this task are provided in Appendix F.1.
+4.3 Time Complexity of HVS
+Method Arch. Lin.
+DINO ViT-S/16 74.06
++ HVS ViT-S/16 74.67
+Impr. +0.61
+SimSiam RN50 68.46
++ HVS RN50 68.98
+Impr. +0.52
+SimCLR RN50 64.31
++ HVS RN50 65.33
+Impr. +1.02
+Table 4: Accounting for computational overhead with 50% resolution (100 epochs).
+The additional forward passes that HVS introduces for the selection phase increase the time complexity of the individual baseline methods.
+Several possible approaches exist to mitigate this overhead, one of which is to reduce the image resolution for the selection step (while still using the original resolution for learning).
+We measured the overhead factors under varying image resolutions for SimSiam and DINO and report them in Table 14 in the appendix.
+Given the 20% (SimSiam) to 30% (DINO) faster training times per batch when using a 50% image reduction, we evaluated this approach and observed that final performances are retained across all tested methods.
+Table 4 provides insights into the HVS’ performance when computational overhead is accounted for in the 50% reduction case.
+Overall, the results validate the effectiveness of HVS in reducing computational demands without compromising on performance.
+We refer to Section H in the appendix for a discussion on further optimizing HVS.
 5 Empirical Analysis of Hard View Selection
@@
 5.1 Q1: Which Patterns Underlying the Hard View Selection Can be Observed?
-When visually studying some of our examples and the views selected by HVS in Figures 2 and 6 (in the appendix), we notice that both geometric and appearance characteristics seem to be exploited.
-E.g., consider the brightness difference between the views of the first two rows in Fig.
+When visually studying examples and the views selected by HVS in Figures 2 and 5 (in the appendix), we notice that both geometric and appearance characteristics seem to be exploited, for instance, see the brightness difference between the views of the first two rows in Fig.
 2.
-Additionally, we observe a generally higher training loss (Fig.
+We also see a generally higher training loss (Fig.
 6 in the appendix) indicative of an increased task difficulty.
-Another key observation is that our adversarial selection chooses view pairs that incorporate zooming in and out or an increased distance between the views (last row in Fig.
-2).
 Logging Augmentation Data
-To assess these observations quantitatively, we logged relevant hyperparameter data during SimSiam training with HVS and conducted multiple studies which we will describe below.
-The data we logged include for each view the sampled parameters from the geometric and appearance data augmentation operations (such as the height/width and coordinates of the crops) as well as color distortion parameters (such as brightness and contrast, see Section D in the appendix for more details), as well as the loss and whether the view was selected or not.
-As evaluated metrics we chose the Intersection over Union (IoU), Relative Distance (image-wise normalized distance of the center points of view pairs), color distortion distance (the Euclidean distance between all four color distortion parameters), and the individual color distortion parameters.
+To assess these observations, we logged relevant hyperparameter data during SimSiam training with HVS.
+The logs include for each view the sampled geometric and appearance parameters from the data augmentation operations (such as the height/width of the crops or the brightness; see Section D in the appendix for more details), as well as the loss and whether the view was selected.
+As evaluated metrics we chose the Intersection over Union (IoU), Relative Distance (normalized distance of the center points views), color distortion distance (the Euclidean distance between all four color distortion parameters), and the individual color distortion parameters.
 Figure 3: Left: In over 40% of the cases, the adversarially selected view pair is also the one with the lowest IoU throughout SimSiam+HVS pretraining.
@@
 4 (left).
-Here, we notice the average IoU value of selected view pairs increases slightly with training, possibly as a reaction of the pretrained model’s embedding representation to the hard view selection.
-Another pattern we observe is that the IoU value with HVS (Fig.
-4 (left) in blue) is smaller and also less constant during training when compared against training without HVS (green).
-We believe this is due to the sample-wise and stateful nature of the adversarial selection as HVS chooses different IoU values from sample to sample and training state to training state.
+An observable pattern is that the IoU value with HVS (Fig.
+4 (left) in blue) is smaller and varies more when compared against training without HVS (green).
+We believe this is due to the sample-wise and stateful nature of the adversarial selection as HVS chooses different IoU values between varying samples and model states.
 Lastly, we assessed the effect of the color augmentation on the pair selection.
-For this study, we sampled one set of color augmentations (as opposed to multiple, i.e.
-one each view) per iteration and applied it to all views.
-Only after identifying the hardest pair, we apply sampled data augmentations for each view individually.
-Consequently, the linear evaluation performance dropped by 0.3% on average.
-Interestingly, as we show in Fig.
+For this study, we sampled one set of color augmentations (as opposed to one for each view) per iteration and applied it to all views.
+Only after identifying the hardest pair, we apply sampled data augmentations to each view.
+As we show in Fig.
 4 (right), the fraction of selected pairs that are also the hardest pairs slightly increases in this case.
 One possible explanation for this is that it reflects the non-negligible role of color variation between views (as shown previously with the importance analysis), where HVS is given less leverage to increase hardness through a static appearance and instead, depends more on leveraging the IoU.
+Another key observation is that HVS often chooses view pairs that incorporate zooming in and out or an increased distance between the views (see last row of Fig.
+2).
+We believe the former relates to the analogy of local-to-global view correspondence described in SimCLR and the DINO papers.
 5.2 Q2: Can a Manual Augmentation Policy be Inferred?
@@
 B+range(0.3-0.35) -0.80 -1.47
-B+range(0.3-0.35)+alternating +0.10 -0.45
+B+range(0.3-0.35)+alt. +0.10 -0.45
 B+range(0.4-0.6) +0.55 -0.40
-B+range(0.4-0.6)+alternating +0.25 -0.20
+B+range(0.4-0.6)+alt. +0.25 -0.20
 B+range(0.1-0.8) -33.95 -1.50
@@
 Table 5: Top-1 lin. eval. accuracies for the manual IoU policy (averaged over 2 seeds).
-Since harder pretraining tasks seem beneficial, and based on the observations made in Q1, a natural question arises: is it possible to mimic the adversarial selection with a manually scripted augmentation policy?
-Such a policy, if effective, would replace HVS and lower the computational cost by avoiding the additional forward passes.
-Since the IoU seems to play an important role, below, we study several possible ways for constructing a simple manual augmentation policy based on IoU.
+Since harder pretraining tasks seem beneficial according to observations made in Q1, a natural question arises: is it possible to mimic the adversarial selection with a manually scripted augmentation policy?
+Such a policy would replace HVS and lower the computational cost .
+Since the IoU plays an important role, below, we study several possible ways for constructing a simple manual augmentation policy based on IoU.
 Deriving an Augmentation Policy
-We implemented the following algorithm based on rejection sampling in the augmentation pipeline: we linearly approximate the IoU value curve from Fig.
-4 (left; in blue) with a start value of 0.30 and end value 0.35 (ignoring the dip in the early phase) to receive an upper bound for the IoU in each training iteration.
-For each iteration, we then check if the pairs exceed the IoU upper bound value and if so, we reject the pair and re-sample a new pair (max.
-512 times).
-This ensures that only view pairs are sampled that employ at least a certain task difficulty (by means of a small enough IoU).
-We varied different hyperparameters (e.g., IoU start/end ranges; inverse schedules; alternating between the IoU schedule and the standard augmentation every other iteration; as well as warmup, i.e.
-no IoU policy for the first 10 epochs).
-Training both SimSiam (ResNet-50) and DINO (ViT-S/16; applied to all global and local heads) models for 100 epochs yielded performance decreases or insignificant improvements, as can be seen in Table 4.
-Moreover, we see an increased risk of model collapse when setting start values too small which can be interpreted as a too hard task.
+We implemented the following rejection sampling algorithm in the augmentation pipeline: we linearly approximate the IoU values from Fig.
+4 (left; in blue) with start and end values of 0.30 and 0.35 (ignoring the dip in the early phase).
+For each iteration, we then check if the pairs exceed the IoU value and if so, we reject the pair and re-sample a new pair.
+This ensures that only pairs are sampled that entail a minimum task difficulty (by means of a small enough IoU).
+We varied different hyperparameters, e.g., IoU start/end ranges, inverse schedules, and alternating between the IoU schedule and the standard augmentation .
+Training both SimSiam and DINO models for 100 epochs yielded performance decreases or insignificant improvements (see Table 5).
 These results indicate that developing a manual policy based on metrics in pixel space such as the IoU is non-trivial.
-In contrast, we believe that HVS is effective since it 1) operates on a similarity level of latent embeddings that may be decorrelated from the pixel space and 2) has access to the current model state.
+Additionally, these results show that transferring such a policy from SimSiam to DINO does not work well, possibly due to additional variations in the augmentation pipeline such as multi-crop.
+In contrast, we believe that HVS is effective and transfers well since it 1) operates on a similarity level of latent embeddings that may be decorrelated from the pixel space and 2) has access to the current model state.
 Figure 4: Left: The average IoU of view pairs selected by SimSiam+HVS (blue) compared against the default SimSiam training (green).
@@
 Assessing the Difficulty of Predicting the Pair Selected by HVS
-To further validate the previous result and to ascertain an upper limit on the performance for predicting the hardest pair, we conducted a second experiment.
-We fitted a gradient boosting classification tree (Friedman, 2001) to predict the selected view pair conditioned on all SimSiam hyperparameter log data from Q1 (see above) except for the flag that indicates whether a view was selected.
+We further validated the previous result and assess an upper limit on the performance for predicting the hardest pair.
+By fitting a gradient boosting classification tree (Friedman, 2001), we predict the selected view pair conditioned on all SimSiam hyperparameter log data from Q1 except for the flag that indicates whether a view was selected.
 As training and test data, we used the logs from two seeds (each 300k samples) and the logs from a third seed, respectively.
@@
 However, the resulting average test performance in all scenarios never exceeded 40%, indicating that it is indeed challenging to predict the hardness of views based on parameter-level data.
-Consequently, we believe these results support the presumption that deriving a policy for controlling and increasing hardness based on geometric and appearance parameters is non-trivial.
-Furthermore, our analysis suggests that such a policy must function on a per-sample basis and have access to the current model state which is the case when applying HVS.
+This further supports our hypothesis that deriving a policy for controlling and increasing hardness based on geometric and appearance parameters is non-trivial and that such a policy must function on a per-sample basis and have access to the current model state as in HVS.
 5.3 Q3: What are the Effects of Empowering the Adversary?
@@
 We optimize a Spatial Transformer Network (STN) (Jaderberg et al., 2015) jointly with the DINO objective and a ViT-tiny.
-Using the inverse of the DINO loss, we train the STN to output the geometric transformation matrices (each 6 parameters) for generating views given an input image that are used for standard CL training with DINO.
+Using the inverse of the DINO loss, we train the STN to output the geometric transformation matrices (each 6 parameters) for generating views .
 We employ and test various transformation types (e.g., rotation, translation, etc.) that influence the capacity of the adversary.
@@
 [^1] For completeness, we also experimented with a cooperative HVS (see Appendix F.2 for details).
-Figure 5: Comparing DINO with DINO+HVS on CIFAR100 for longer trainings.
-Longer Training
-Despite the inherent overhead associated with additional forward passes of HVS, our results clearly show that using HVS is also beneficial for long trainings (see 300-epoch improvements in Table 1).
-To delve deeper into the behavior of longer trainings under our computational constraints, we conduct several training runs with progressively higher epochs on the CIFAR100 dataset, using the same initial learning rate and a cosine schedule.
-The comparative analysis in Fig.
-5 shows, except for some overfitting, no evidence of diminishing returns as the improvement remains consistent throughout training nor signs of instability (e.g., collapse) due to hard tasks.
 6 Conclusion
-In this study, we presented HVS, a new data augmentation and learning strategy for contrastive learning designed to challenge pretrained models with more demanding samples.
-This method, while straightforward in its design, has proven to be a powerful tool, pushing the boundaries of the traditional random view generation in contrastive learning.
-When combined with methods like DINO, SimSiam, and SimCLR, HVS consistently showcased improvements in linear evaluation and a diverse set of transfer tasks.
+In this study, we presented HVS, a new data augmentation and learning strategy for Self-Supervised Learning designed to challenge pretrained models with more demanding samples.
+This method, while straightforward in its design, has proven to be a powerful tool, pushing the boundaries of the traditional random view generation in SSL.
+When combined with methods like DINO, SimSiam, iBOT, and SimCLR, HVS consistently showcased improvements in linear evaluation and a diverse set of transfer tasks.
 As the landscape of self-supervised learning continues to evolve, methods like HVS underscore the potential for further advancements in the field.
 By continuously challenging models with harder samples and refining augmentation techniques, we anticipate even greater strides in model performance and generalization in future research endeavors.
-While HVS incurs a computational overhead due to its additional forward passes, our preliminary experiments indicate no signs of diminishing returns with extended training durations.
+While HVS introduces computational overhead due to its additional forward passes, our findings demonstrate that this overhead is significantly mitigated when using small resolution images in HVS without impeding its performance.
 This suggests that HVS is likely to continue to outperform baselines regardless of the training length.
-Looking ahead, we believe there is ample room for refining and expanding HVS, such as further investigating the number of sampled views based on the current learning state, enhancing efficiency through checkpointing activations, bypassing forwarding of similar pairs, or augmentation distributions that allow controlling task difficulty.
 Appendix A Examples Sampled by HVS
@@
 As can be seen in Fig.
-9, the metric with the highest predictive capacity on the loss is the IoU with an importance of 15.2% followed by brightness with 5.1%.
+8, the metric with the highest predictive capacity on the loss is the IoU with an importance of 15.2% followed by brightness with 5.1%.
 The relative distance has an importance of 3.3%, the colorjitter distance 2.3%, the contrast 1.6%, the saturation 1.4%, hue 0.6%, and all parameters jointly 1.7%.
@@
 DINO is trained with its original contrastive objective, where the STN is trained by inverting the gradient after the DINO during backpropagation.
-{tikzpicture}
-[…]
-[…]
-[…]
+InputSTNCropsViewsDINOLossaugm.maximizeminimizePenalty
 Figure 9: Illustration of adversarial learning with a Spatial Transformer Network (STN) jointly with contrastive learning (here: DINO).
@@
 Table 13: Hyperparameters for VOC object detection and COCO instance segmentation.
-Appendix H Hard View Selection Objectives
-H.1 SimCLR
+Appendix H Computational Overhead of HVS
+In Table 14, we compare the computational overhead factors across multiple baselines.
+The details of hardware and software used for this analysis are: NVIDIA RTX 3080, AMD R7 5800X, 32GB RAM, Ubuntu 22.04, PyTorch 2.0.1, CUDA 11.8.
+For DINO’s 2 global and 8 local views (default), applying HVS with nviews=2 sampled for each original view results in 4 global and 16 local views.
+Since considering all combinations would yield over 77k unique comparisons (4 over 2 times 16 over 8), to remain tractable, we limit the number of total comparisons to 128.
+Overall, reducing the image resolution can significantly reduce the time complexity of HVS.
+For instance, in the case of DINO ViT-S/16, the reduction leads to approximately 30% faster training times and as reported in the main paper, a 50% reduction allows retaining original performance.
+Res. SimSiam (RN50) DINO (ViT)
+100% x1.55 x2.15
+75% x1.45 x1.68
+50% x1.24 x1.56
+25% x1.16 x1.5
+Table 14: Time slowdown factors under different image resolutions for the selection in HVS.
+While technically there can be a memory overhead with HVS, with the number of sampled views chosen in this paper, the backward pass of the methods that compute gradients only for the selected view pair still consumes more memory than the selection part of HVS (even for 8 sampled views in SimSiam).
+Note, that selection and the backward computation are never executed at the same time but sequentially.
+We emphasize that the time overhead factors were measured without any optimization of HVS’ efficiency and, in our view, there are multiple ways to improve it.
+In our work, we opted for the easiest implementation possible to showcase that selecting harder views dependent on the model learning state can help boost performance in contrastive learning.
+Going forward with more compute-efficient HVS solutions, one could think of:
+  • using embeddings of views from “earlier” layers in the networks or
+  • using 4/8 bit low-precision for the view selection or
+  • using one GPU just for creating embeddings and selecting the hardest views while the remaining GPUs are used for learning or
+  • switching between HVS and the standard pipepline in alternating fashion or
+  • bypassing forwarding of similar pairs and more
+Appendix I Hard View Selection Objectives
+I.1 SimCLR
 In this section, we are going to introduce the application of HVS with the SimCLR objective.
@@
 and $k,l\in\{1,2,\ldots,N\}$ and $i,j\in\{1,2,\ldots,M\}$.
-In this formulation, we iterate over all possible pairs of augmented images derived from the same original image across all ”N” augmentations, and compute the loss for each pair.
-The pair that gives the highest loss is then used in the optimization process.
-It is important to note that the temperature parameter $\tau$ and the similarity function remain as defined in the original loss function.
+For each iteration, we evaluate all possible view pairs and contrast each view against every other example in the mini-batch.
+Intuitively, the pair that yields the highest loss is selected, which is the pair that at the same time minimizes the numerator and maximizes the denominator in the above equation.
+In other words, the hardest pair is the one, that has the lowest similarity with another augmented view of itself and the lowest dissimilarity with all other examples.
+Appendix J Attention Maps
+In this section, we visualize attention maps as a way to qualitatively study the potential effects and differences of the HVS learning method and their resulting features.
+The provided attention maps are from a DINO ViT-S/16 100 model, and we contrast the attention for each HVS and the baseline.
+All input images are from the ImageNet-1k validation split.
+The color code used depicts strong attention in yellow and weak attention in green and attentions from the HVS model are shown in the top row and attentions from the baseline in the bottom row, respectively.
+To summrize this study, we do not see apparent strong differences between HVS and the baseline.
+What we occasionally notice is that the HVS models seem to capture the shape of some subtle/indistinct objects better (e.g.
+the creek/river in the valley in Fig.
+12, the lizard in Fig.
+13, or the speaker in Fig.
+14 or focus slightly more on the context and background (Fig.
+16, 17, and 18).
+Figure 11:
+Figure 12:
+Figure 13:
+Figure 14:
+Figure 15:
+Figure 16:
+Figure 17:
+Figure 18:
+Figure 19:

### v6

@@
-Hard View Selection for Self-Supervised Learning
+Beyond Random Augmentations: Pretraining with Hard Views
 Abstract
-Many Self-Supervised Learning (SSL) methods train their models to be invariant to different “views” of an image and considerable efforts were directed towards improving pre-text tasks, architectures, or robustness.
-However, most SSL methods remain reliant on the random sampling of operations within the image augmentation pipeline, such as the random resized crop operation.
-We argue that the role of the view generation and its effect on performance has so far received insufficient attention.
-To address this, we propose an easy, learning-free, yet powerful Hard View Selection (HVS) strategy designed to extend the random view generation to expose the pretrained model to harder samples during SSL training.
-It encompasses the following iterative steps: 1) randomly sample multiple views and create pairs of two views, 2) run forward passes for each view pair on the currently trained model, 3) adversarially select the pair yielding the worst loss depending on the current model state, and 4) run the backward pass with the selected pair.
-As a result, HVS consistently achieves accuracy improvements between 0.91% and 1.93% on ImageNet linear evaluation and similar improvements on transfer tasks across DINO, SimSiam, iBOT and SimCLR.
-We provide studies to shed light on the inner workings and show that, by naively using smaller resolution images for the selection step, we can significantly reduce the computational overhead while retaining performance.
-Surprisingly, even when accounting for the computational overhead incurred by HVS, we achieve performance gains between 0.52% and 1.02% and closely rival the 800-epoch DINO pretraining with only 300 epochs.
+Self-Supervised Learning (SSL) methods typically rely on random image augmentations, or views, to make models invariant to different transformations.
+We hypothesize that the efficacy of pretraining pipelines based on conventional random view sampling can be enhanced by explicitly selecting views that benefit the learning progress.
+A simple yet effective approach is to select hard views that yield a higher loss.
+In this paper, we propose Hard View Pretraining (HVP), a learning-free strategy that extends random view generation by exposing models to more challenging samples during SSL pretraining.
+HVP encompasses the following iterative steps: 1) randomly sample multiple views and forward each view through the pretrained model, 2) create pairs of two views and compute their loss, 3) adversarially select the pair yielding the highest loss according to the current model state, and 4) perform a backward pass with the selected pair.
+In contrast to existing hard view literature, we are the first to demonstrate hard view pretraining’s effectiveness at scale, particularly training on the full ImageNet-1k dataset, and evaluating across multiple SSL methods, ConvNets, and ViTs.
+As a result, HVP sets a new state-of-the-art on DINO ViT-B/16, reaching 78.8% linear evaluation accuracy (a 0.6% improvement) and consistent gains of 1% for both 100 and 300 epoch pretraining, with similar improvements across transfer tasks in DINO, SimSiam, iBOT, and SimCLR.
 1 Introduction
-Figure 1: HVS first samples $N$ views, pairs them, and adversarially selects the hardest pair, i.e., the one with the worst loss according to the current model state.
-Various approaches to learn effective and generalizable visual representations in Self-Supervised Learning (SSL) exist.
-One way to categorize many SSL methods is to distinguish generative and discriminative approaches.
-While generative methods aim at generating image input, discriminative methods, with contrastive learning (Hadsell et al., 2006; He et al., 2020) being the most renowned approach, aim at learning a latent representation in which similar image views are closely and dissimilar ones are distantly located.
+Learning effective and generalizable visual representations in Self-Supervised Learning (SSL) has been approached in various ways.
+Many SSL methods can be broadly categorized into generative and discriminative approaches (Chen et al., 2020a).
+Generative methods focus on generating image inputs, while discriminative methods, particularly contrastive learning (Hadsell et al., 2006; He et al., 2020), aim at learning latent representations in which similar image views are located closely, and dissimilar ones distantly.
 Such views are generated by applying a sequence of (randomly sampled) image transformations and are usually composed of geometric (cropping, rotation, etc.) and appearance (color distortion, blurring, etc.) transformations.
-A body of literature (Chen et al., 2020a; Wu et al., 2020; Purushwalkam & Gupta, 2020; Wagner et al., 2022; Tian et al., 2020b) has illuminated the effects of image views on representation learning and identified random resized crop (RRC) transformation, which randomly crops the image and resizes it back to a fixed size, as well as color distortion as one of the most essential ones for effective contrastive learning.
-However, despite this finding and to our best knowledge, little research has gone into identifying more effective ways for selecting or generating views to improve performance.
-In this paper, we seek to extend the random view generation to expose the pretrained model to harder samples during CL training.
-Figure 2: Examples (left) and sampled views (right) after resizing and appearance transformations. Pairs selected by HVS are shown with a solid frame.
-Supervised Learning works that address this shortcoming and which attempt to leverage task difficulty include Spatial Transformer Networks (Jaderberg et al., 2015) and Adversarial AutoAugment (Zhang et al., 2020), which either directly learn geometric transformations or generate augmentation policies that yield hard samples.
-However, these approaches have not been used successfully for unsupervised learning and Adversarial AutoAugment does not include RRC in its search space.
-In contrast to what we propose, integrating any of these methods into existing self-supervised learning pipelines is non-trivial.
-Related SSL approaches exist that aim at controlling view hardness through adversarial Shi et al.
-(2022); Tamkin et al.
-(2021) or cooperative Hou et al.
-(2023) models.
-A work that takes a similar perspective like us in leveraging the information content of views for SSL is (Tian et al., 2020b).
-They discover sweet spots for data augmentation hyperparameters that are widely used and adversarially learn flow-based models that generate novel color space-views which are then used for contrastive learning.
-While these works offer important insights into view optimality, they usually require learning additional models prior to the actual SSL training or non-trivial changes to the pipeline and thus, are not easily integrated into existing SSL methods.
-In contrast, we propose a fully learning-free and easy-to-integrate approach.
-Importantly, and different to Tian et al.
-(2020b), our approach leverages the current model state and operates on a per-sample basis during pretraining.
-It applies the following simple strategy: we sample a few views and forward each pair through the currently trained model.
-Next, we identify the pair that yields the worst loss among all pairs, use it for the backward pass, and repeat (see Fig.
-1).
-While being agnostic to any contrastive and non-contrastive loss, this strategy exposes the pretrained model to samples that the model at any given point in the training trajectory finds challenging.
-As our results show, sampling only four views are sufficient to achieve noticeable improvements.
+Prior work (Chen et al., 2020a; Wu et al., 2020; Purushwalkam & Gupta, 2020; Wagner et al., 2022; Tian et al., 2020b) has identified random resized crop (RRC), which randomly crops the image and resizes it back to a fixed size, as well as color distortion as critical transformations for effective representation learning.
+However, despite this finding and to our best knowledge, little research has gone into identifying more effective ways for generating views to improve performance.
+Figure 1: (a) HVP first samples $N$ views, pairs them, and adversarially selects the hardest pair, i.e., the one with the worst loss according to the current model state.
+(b) Examples (left) and sampled views (right) after transformations.
+Hard pairs selected by HVP are shown with a solid frame.
+Existing SSL approaches that attempt to control the hardness of views include adversarial (Shi et al., 2022; Tamkin et al., 2021) or cooperative (Hou et al., 2023) techniques.
+For instance, Tian et al.
+(2020b) use mutual information theory to adversarially learn view generators.
+These methods offer valuable insights into optimizing views but often introduce additional complexity, such as requiring additional models or significant changes to the SSL pipeline, limiting their practicality in state-of-the-art models where resource constraints are already a concern.
+Similar to what we propose, Koçyigit et al.
+(2023) introduce a hard view sampling strategy but targeting the acceleration of pretraining.
+Moreover, their approach requires tuning hyperparameters like learning rate and augmentation magnitude, which can introduce confounding variables.
+Furthermore, neither Koçyigit et al.
+(2023) nor Tian et al.
+(2020b) validate their methods on larger datasets such as ImageNet-1k (Deng et al., 2009), or across larger model architectures, limiting their scalability and applicability.
+Building on these observations, we propose Hard View Pretraining (HVP), a fully learning-free, easy-to-integrate approach designed to improve standard pretraining methods without the need for additional model training or complex modifications.
+Our method leverages the current model state to select challenging samples during pretraining by adversarially sampling pairs of views and selecting the pair that yields the highest loss according to the model‘s current state (see Fig.
+1(a)).
+Unlike previous approaches, HVP requires no hyperparameter tuning and demonstrates scalability to large datasets like ImageNet-1k, offering an efficient and practical solution for improving SSL pretraining.
+To the best of our knowledge, we are the first to demonstrate the effectiveness of a hard view sampling strategy at scale, particularly on modern architectures like Vision Transformers (ViTs) and training on the full ImageNet dataset.
+Our approach not only integrates seamlessly with recent state-of-the-art SSL methods but also showcases consistent improvements across both convolutional architectures and ViTs, validating its robustness and scalability.
 Overall, our contributions can be summarized as follows:
-  • We propose Hard View Selection (HVS), a new learning method complementing SSL that is easy to integrate and extends the broadly-used random view generation to automatically expose the model to harder samples during pretraining.
-The only requirement of HVS is that it needs to be able to compute sample-wise losses;
-  • We demonstrate the effectiveness and compatibility of our approach using ImageNet (Deng et al., 2009) pretraining across popular and widely-adopted SSL methods that cover a diverse range of objectives such as SimSiam (Chen & He, 2021), DINO (Caron et al., 2021), iBOT (Zhou et al., 2021), and SimCLR (Chen et al., 2020a);
-  • HVS improves all baselines noticeably compared to using the common view generation, and we show this on a diverse set of transfer datasets and tasks, including linear evaluation, finetuning, object detection and segmentation;
-  • We present insights into the underlying mechanisms of HVS such as the effect of changing the number of sampled views or color augmentation.
-We make our PyTorch (Paszke et al., 2019) code, models, and all used hyperparameters publicly available under https://anonymous.4open.science/r/hard-view-selection/.
+  • We propose Hard View Pretraining (HVP), an easy-to-use method complementing SSL by extending the common random view generation to automatically expose the model to harder samples during pretraining.
+HVP simply requires the ability to compute sample-wise losses;
+  • We demonstrate the effectiveness and compatibility of our approach using ImageNet-1k pretraining across four popular SSL methods that cover a diverse range of discriminative objectives such as SimSiam (Chen & He, 2021), DINO (Caron et al., 2021), iBOT (Zhou et al., 2021), and SimCLR (Chen et al., 2020a);
+  • HVP achieves a new state-of-the-art result on DINO ViT-B/16, improving over the officially reported baseline of 78.2% linear evaluation accuracy by reaching 78.8% (400 epochs).
+HVP also consistently improves all other baselines by an average of 1% in linear evaluation on ImageNet across 100 and 300 epoch-pretraining runs;
+  • We show similar improvements on a diverse set of transfer tasks, including finetuning, object detection, and segmentation, and present insights into the underlying mechanisms and robustness of HVP.
+We make our PyTorch Paszke et al. (2019) code, models, and all used hyperparameters publicly available under https://github.com/automl/hvp.
 2 Related Work
-2.1 Contrastive and Non-Contrastive Learning in SSL
-The core idea behind the contrastive learning framework is to learn image representations by contrasting positive pairs (two views of the same image) against negative pairs (two views of different images) (Hadsell et al., 2006).
+2.1 Discriminative Self-Supervised Learning
+The core idea behind the discriminative learning framework (Chen et al., 2020a) is to learn image representations by contrasting positive pairs (two views of the same image) against negative pairs (two views of different images) (Hadsell et al., 2006).
 To work well in practice and to prevent model collapse, contrastive learning methods often require a large number of negative samples (Wu et al., 2018; van den Oord et al., 2018; Chen et al., 2020a; He et al., 2020; Tian et al., 2020a; Chen et al., 2020b) stored in memory banks (Wu et al., 2018; He et al., 2020) or, for instance, in the case of SimCLR, implicitly in large batches (Chen et al., 2020a).
-Non-contrastive approaches, such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), DINO (Caron et al., 2021) and others (Zbontar et al., 2021; Caron et al., 2020; Ermolov et al., 2021), are able to only use positive pairs without causing model collapse but rely on other techniques, such as Siamese architectures, whitening of embeddings, clustering, maximizing the entropy of the embeddings, momentum encoders, and more.
-2.2 Optimizing for Hard Augmentations and Views in SSL
-Past work has shown that optimizing augmentation policies or views directly can be beneficial.
-Hence, the realm of learning task-specific augmentation policies based on data has seen quick development (Cubuk et al., 2019; Ho et al., 2019; Lin et al., 2019; Zhang et al., 2020; Hataya et al., 2020; Müller & Hutter, 2021).
-Most related to our work among these approaches is Adversarial AutoAugment (AAA) (Zhang et al., 2020) that proposes to generate augmentation policies yielding hard samples by learning with an adversarial supervised objective.
-In a similar, but cooperative setting, MADAug (Hou et al., 2023) learns policies yielding progressively harder samples.
-In contrast to our work, however, AAA and MADAug do not include the random resize crop operation into their search spaces and thus, are limited in leveraging the geometric perspective to control view hardness.
-Unlike learning hard augmentations, approaches exist that optimize directly on the pixel level to control view hardness.
-For instance, Shi et al.
-(2022) adversarially learn semantically meaningful masks for a pretrained network to inpaint.
-Another approach perturbs input images using an adversarial model (Tamkin et al., 2021).
-Tian et al.
-(2020b) leverage, like us, the view content for better performance.
-Under the lens of mutual information (MI), they derive an objective to adversarially learn flow-based view generators for contrastive learning.
-While these methods can complement our approach, they are less easily integrated into existing contrastive pipelines due to the need for adversarial network training.
-Moreover, Tian et al.
-(2020b) does not use the current model state to control learning hardness and, while their studies incorporate distance metrics, they do not consider the Intersection over Union (IoU) metric which we extensively make use of.
+Non-contrastive approaches, such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), DINO (Caron et al., 2021) and others (Zbontar et al., 2021; Caron et al., 2020; Ermolov et al., 2021), can only use positive pairs without causing model collapse but rely on other techniques, such as Siamese architectures, whitening of embeddings, clustering, maximizing the entropy of the embeddings, momentum encoders, and more.
+2.2 Optimizing for Hard Views in SSL
+Due to its performance-improving benefits, the realm of learning task-specific augmentation policies based on data has seen quick development (Cubuk et al., 2019; Ho et al., 2019; Lin et al., 2019; Zhang et al., 2020; Hataya et al., 2020; Hou et al., 2023; Müller & Hutter, 2021).
+However, these approaches do not include the random resize crop operation in their search spaces, limiting the control of view hardness.
+Similar to us, Koçyigit et al.
+(2023) uses the current model state for selecting hard views.
+However, their approach requires controlling learning hyperparameters, while mostly training on a smaller version of ImageNet and ResNets (He et al., 2016) only.
+We offer a more complete analysis of hard view pretraining, demonstrating the benefits without potential confounding factors such as hyperparameter adjustments.
+We also focus on performance rather than pretraining speedups and employ higher budgets (longer pretraining, larger batch sizes) on the full ImageNet dataset and both ResNets and ViTs (Dosovitskiy et al., 2020)).
+Other works utilize additional networks for view generation, such as Tamkin et al.
+(2021); Shi et al.
+(2022); Tian et al.
+(2020b) (adversarial view generators), Peng et al.
+(2022) (localization network for semantic awareness), Li et al.
+(2024) (pretrained generative models to enhance augmentation quality), and Han et al.
+(2023) (leveraging a pretrained GAN in a SimCLR-only setting).
+However, unlike HVP, all these methods add non-trivial complexity to the training pipeline by requiring learning auxiliary and adversarial components or are often limited in their applicability across different SSL frameworks (e.g., by requiring negative view pairs).
 3 Method
-3.1 Contrastive Learning Framework
-In this section, we introduce our approach which is also depicted in Algorithm 12.
-Many different contrastive learning (He et al., 2020) objectives exist, each characterized by variations stemming from design choices, such as by the use of positive and negative samples or asymmetry in the encoder-projector network structure.
-For simplicity of exposure, we will introduce our approach based on the SimSiam objective Chen & He (2021), but we do note that our method can be used with any other discriminative SSL objective that allows the computation of a sample-wise losses.
+3.1 Self-supervised Learning Framework
+In this section, we introduce our approach, which is also depicted in Algorithm 12.
+Many different self-supervised discriminative learning (He et al., 2020) objectives exist, each characterized by variations stemming from design choices, such as by the use of positive and negative samples or asymmetry in the encoder-projector network structure.
+For simplicity of exposure, we will introduce our approach based on the SimSiam objective (Chen & He, 2021), but we do note that our method can be used with any other discriminative SSL objective that allows the computation of sample-wise losses.
 SimSiam works as follows.
@@
 where D denotes the negative cosine similarity function. Intuitively, when optimizing $\theta$, the embeddings of the two augmented views are attracted to each other.
-3.2 Hard View Selection
-As previously described, Hard View Selection extends the random view generation by sampling adversarially harder views during pretraining.
+3.2 Pretraining with Hard Views
+We now formalize how we expose the model to more challenging views during pretraining.
+In a nutshell, Hard View Pretraining extends the random view generation by sampling adversarially harder views during pretraining.
 Instead of having two sets of augmentations $A$ and $B$, we now sample $N$ sets of augmentations, denoted as $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{N}\}$.
@@
 We then define a new objective function that seeks to find the pair $(x_{i}^{A_{k}},x_{i}^{A_{l}})$ of a given image $x_{i}$ that yields the highest loss:
-$(x_{i}^{A_{k}},x_{i}^{A_{l}})=\operatorname*{arg\,max}_{k,l;k\neq l}\mathcal{L}(\theta)_{i,k,l}=\operatorname*{arg\,max}_{k,l;k\neq l}\frac{1}{2}\left(D(z_{i}^{A_{k}},h_{i}^{A_{l}})+D(z_{i}^{A_{l}},h_{i}^{A_{k}})\right),$ (2)
-where $\mathcal{L}(\theta)_{i,k,l}$ simply denotes a sample-wise variant of Eq.
-1.
-Overall, we first generate $N$ augmented views for each image $x_{i}$ in the minibatch.
-Then, we create all combinatorially possible $\binom{N}{2}$ pairs of augmented images for $x_{i}$.
-Then, we use Eq.
-2 to compute the sample-wise loss for each pair which corresponds to invoking forward passes.
-We then select all pairs that yielded the highest loss to form the new hard minibatch of augmented sets $\textbf{x}^{A_{k_{*}}}$ and $\textbf{x}^{B_{l_{*}}}$, discard the other pairs and use the hard minibatch for optimization.
+$(x_{i}^{A_{k}},x_{i}^{A_{l}})$ $=\operatorname*{arg\,max}_{k,l;k\neq l}\mathcal{L}(\theta)_{i,k,l}$ (2)
+$=\operatorname*{arg\,max}_{k,l;k\neq l}\frac{1}{2}\left(D(z_{i}^{A_{k}},h_{i}^{A_{l}})+D(z_{i}^{A_{l}},h_{i}^{A_{k}})\right),$
+where $\mathcal{L}(\theta)_{i,k,l}$ simply denotes a sample-wise variant of Eq. 1.
 1:Input: Number of views $N\geq 2$, batch size $M$,
-2:augmentation distribution $\mathcal{T}$, encoder $f_{\theta}$, projector $g_{\theta}$
-3:for each $x_{i}$ in the sampled minibatch $\left\{x_{i}\right\}_{i=1}^{M}$ do
+2:augmentation distribution $\mathcal{T}$, model $f$
+3:for each $x_{i}$ in the sampled batch $\left\{x_{i}\right\}_{i=0}^{M}$ do
 4:     Sample $N$ augmentations: $A=\left\{t_{n}\sim\mathcal{T}\right\}_{n=0}^{N}$
 5:     Create augmented views: $\textbf{x}_{i}^{A}=\left\{t_{n}(x_{i})\right\}_{n=0}^{N}$
-6:     Forward all views through $f_{\theta}$ and $g_{\theta}$ to get $z$ and $h$
-7:     Create all $\binom{N}{2}$ possible pairs $\textbf{x}_{i}^{A_{k}}\times\textbf{x}_{i}^{A_{l}}$, $k\neq l$
-8:     Select the hard pair $(x_{i}^{A_{k_{*}}},x_{i}^{A_{l_{*}}})$ maximizing Eq. 2
-9:Create the minibatch $(\textbf{x}^{A_{k_{*}}},\textbf{x}^{A_{l_{*}}})$ consisting of hard pairs
-10:Proceed with standard contrastive learning training
-to update $\theta$ with the hard batch
-11:Repeat for all minibatches
-12:return Optimized $\theta$
-Algorithm 1 Contrastive Learning with HVS
+6:     Forward all views through $f$
+7:     Create all $\binom{N}{2}$ view pairs $\textbf{x}_{i}^{A_{k}}\times\textbf{x}_{i}^{A_{l}}$, $k\neq l$
+8:     Add hard pair $(x_{i}^{A_{k_{*}}},x_{i}^{A_{l_{*}}})$ that maximizes
+9:    Eq. 2 to the new batch with only hard pairs
+10:Proceed with standard SSL training
+11:Repeat for all batches
+12:return Pretrained model $f$
+Algorithm 1 Pretraining with Hard Views
+Overall, we first generate $N$ augmented views for each image $x_{i}$ in the minibatch.
+Then, we forward these augmented views through the networks and create all combinatorially possible $\binom{N}{2}$ pairs of augmented images.
+Subsequently, we use Eq.
+2 to compute the sample-wise loss for each pair.
+We then select all pairs that yielded the highest loss to form the new hard minibatch of augmented sets $\textbf{x}^{A_{k_{*}}}$ and $\textbf{x}^{B_{l_{*}}}$, discard the other pairs and use the hard minibatch for optimization.
 As shown in Algorithm 12, we repeat this process in each training iteration.
-This modification introduces a more challenging learning scenario in which the model is encouraged to learn more discriminative features by being exposed to harder views.
-While the additional forward passes come with a certain overhead (about a factor of 1.55$\times$ for SimSiam), as we will demonstrate in our experiments, the focus on harder views leads to consistent improvements.
-While we exemplified the integration of HVS with the SimSiam objective, integrating it into other contrastive methods is as straightforward.
-The only requirement of HVS is to be able to compute sample-wise losses (in order to select the views with the highest loss).
-In our experiments section and in addition to SimSiam, we study the application of HVS to the objectives of DINO and SimCLR (see also Appendix I.1 for a formal exemplary definition for the integration of HVS into SimCLR).
-3.3 Implementation and Evaluation Protocols
-Implementation
+Intuitively, our approach introduces a more challenging learning scenario in which the model is encouraged to learn more discriminate features by being exposed to harder views.
+In the early stage of training, the embedding space lacks a defined structure for representing similarity among views.
+As training progresses, our method refines the concept of similarity through exposure to views that, from the perspective of the model, remain challenging given its current state.
+By limiting the number of sampled views, we upper-bound the difficulty of learning to prevent tasks from becoming too difficult and hindering learning.
+This ensures a controlled evolution of the embedding space, where the model’s perception of difficulty is continuously challenged in tandem with its growing capacity to differentiate views.
+Consequently, HVP can be seen as a regularization that prevents the model from overfitting to easy views.
+While we exemplified the integration of HVP with the SimSiam objective, integrating it into other contrastive methods is as straightforward.
+The only requirement of HVP is to be able to compute sample-wise losses (to select the views with the highest loss).
+In our experiments section and in addition to SimSiam, we study the application of HVP to the objectives of DINO, iBOT, and SimCLR (see also Appendix K.1 for a formal exemplary definition for the integration of HVP into SimCLR).
+4 Implementation and Evaluation Protocols
+4.1 Implementation
 We now describe the technical details of our approach.
-HVS can be used with any contrastive method that allows computing sample-wise losses, and the only two elements in the pipeline we adapt are: 1) the data loader (needs to sample $N$ views for each image) and 2) the forward pass (invokes the select function to identify and return the hard views).
-The image transformation distribution $\mathcal{T}$ of the baselines is left unchanged.
-Note, for the view selection one could simply resort to RRC only and apply the rest of operations after the hard view selection (see Section 5.1 for a study on the influence of appearance on the selection).
-All experiments were conducted with $N=4$ sampled views, yielding $\binom{N}{2}=6$ pairs to compare, except for DINO that uses 10 views (2 global, 8 local heads) per default.
-For DINO, we apply HVS to both global and local heads but to remain tractable, we upper-bound the number of total pair comparisons to 128.
+HVP can be used with any SSL method that allows computing sample-wise losses, and the only two elements in the pipeline we adapt are: 1) the data loader (which now needs to sample $N$ views for each image) and 2) the forward pass (which now invokes a select function to identify and return the hard views).
+The image transformation distribution $\mathcal{T}$ taken from the baselines is left unchanged.
+Note, for the view selection one could simply resort to random resized crop (RRC) only and apply the rest of the operations after the hard view selection (see Section 6.1 for a study on the influence of appearance on the selection).
+All experiments were conducted with $N=4$ sampled views, yielding $\binom{N}{2}=6$ pairs to compare, except for DINO which uses 10 views (2 global, 8 local heads) per default.
+For DINO, we apply HVP to both global and local heads but to remain tractable, we upper-bound the number of total pair comparisons to 128.
 SimCLR uses positive and negative samples.
-In accordance with the simplicity of HVS, we do not alter its objective, which naturally leads to selecting hard views that are adversarial to positive and “cooperative” to negative views.
+Following the simplicity of HVP, we do not alter its objective, which naturally leads to selecting hard views that are adversarial to positive and “cooperative" to negative views.
+For iBOT, we use the original objective as defined by the authors with global views only.
 Method Arch. 100 epochs 300 epochs
@@
 DINO ViT-S/16 73.52 68.80 75.48 72.62
-+ HVS ViT-S/16 74.67 70.72 76.56 73.64
-Improvement +1.15 +1.92 +1.08 +1.02
++ HVP ViT-S/16 74.67 70.72 76.56 73.65
+Impr. +1.15 +1.92 +1.08 +1.03
 DINO RN50 71.93 66.28 75.25 69.53
-+ HVS RN50 72.87 67.33 75.65 70.05
-Improvement +0.94 +1.05 +0.40 +0.52
++ HVP RN50 72.87 67.33 75.65 70.05
++0.94 +1.05 +0.40 +0.52
 SimSiam RN50 68.20 57.47 70.35 61.40
-+ HVS RN50 68.98 58.97 70.90 62.97
-Improvement +0.78 +1.50 +0.55 +1.57
-SimCLR RN50 63.40 52.83 65.50 55.65
-+ HVS RN50 65.33 54.76 67.20 57.31
-Improvement +1.93 +1.85 +1.70 +1.66
-iBOT ViT-S/16 69.50 62.82 72.48 66.60
-+ HVS ViT-S/16 70.41 63.05 73.99 67.12
-Improvement +0.91 +0.23 +1.51 +0.52
-Table 1: Average top-1 linear and k-NN classification accuracy on the ImageNet validation set for 100 and 300-epoch pretrainings. For SimCLR we run 200 instead of 300 epochs.
-Evaluation Protocols
++ HVP RN50 68.98 58.97 70.90 62.97
++0.78 +1.50 +0.55 +1.57
+SimCLR RN50 63.37 52.83 65.50 55.65
++ HVP RN50 65.33 54.76 67.30 56.80
++1.96 +1.93 +1.80 +1.15
+iBOT ViT-S/16 69.55 62.93 72.76 66.92
++ HVP ViT-S/16 70.27 62.75 73.99 67.16
++0.73 -0.18 +1.23 +0.24
+Table 1: Average top-1 linear and k-NN classification accuracy on the ImageNet validation set for 100 and 300-epoch pretrainings across 3 seeds.
+4.2 Evaluation Protocols
 We now describe the protocols used to evaluate the performance in our main results section.
@@
 Here, we discuss our main results on image classification, object detection, and segmentation tasks.
-All results are self-reproduced using the baseline code and available hyperparameters, and are averaged over 3 seeds (iBOT over 2 seeds).
-All hyperparameters are reported in Appendix G.
+All results are self-reproduced using the original baseline code and hyperparameters (see Appendix Section A for details).
 5.1 Evaluations on ImageNet
@@
 For DINO, we additionally compare ResNet-50 (He et al., 2016) against the ViT-S/16 (Dosovitskiy et al., 2020) architecture.
-Our method compares favourably against all baselines with an increased performance of approximately $1\%$ on average, showing the benefit of sampling hard views.
-We highlight that our 300-epoch DINO+HVS model exhibits strong performance, closely rivaling the officially reported 800-epoch DINO performance of 77%, with a mere ${\sim}0.4\%$ performance gap.
-This achievement remains favorable, even when factoring in the 2$\times$ slowdown (DINO with 2 global, 8 local heads) induced by our method.
-Due to limited compute resources, we run the pretrainings for 100 and 300 epochs (200 epochs for SimCLR) and batch sizes of 512 (100 epoch) or 1024 (200 & 300 epoch trainings), respectively.
-This choice is in line with a strategy that favors the evaluation of a diverse and larger set of baselines over the evaluation of a less diverse and smaller set and underpins the broad applicability of HVS.
-We primarily ran our experiments with 8xNVIDIA GeForce RTX 2080 Ti nodes, with which the cheapest runs required ${\sim}3.5$ days for pretraining and linear evaluation and the most expensive runs required ${\sim}25$ days.
-Moreover, we emphasize that HVS is insensitive to the baseline hyperparameters and simply reusing these consistently resulted in improvements of the reported magnitudes across all experiments.
+We point out that both methods, vanilla, and vanilla+HVP always receive the same number of data samples for training.
+Our method compares favorably against all baselines with an increased performance of approximately $1\%$ on average for 100 and 300 epoch pretraining, showing the benefit of sampling hard views.
+Due to limited computing resources, we run the majority of pretrainings in this paper for 100 epochs and 300 epochs (200 epochs for SimCLR) and batch sizes of 512 (100 epoch) or 1024 (200 & 300 epoch trainings), respectively.
+This choice is in line with a strategy that favors the evaluation of a diverse and larger set of baselines over the evaluation of a less diverse and smaller set and underpins the broad applicability of HVP.
+We primarily ran our experiments with 8xNVIDIA GeForce RTX 2080 Ti nodes, with which the pretraining and linear evaluation duration ranged from ${\sim}3.5$ to ${\sim}25$ days.
+While HVP requires roughly 2x the training time of the DINO baseline (see Appendix J for further discussion on the time complexity of HVP), this investment translates directly into improved performance.
+Rather than prioritizing efficiency, our focus was on achieving state-of-the-art results, highlighting HVP’s flexibility and robustness across training durations.
+To showcase this point, we explored the scalability of HVP with larger models and extended training schedules and achieved a new state-of-the-art result of 78.8% accuracy in linear evaluation, improving over the officially reported baseline of 78.2% on DINO ViT-B/16 (400 epochs).
+For k-NN classification, the same model similarly surpassed the DINO baseline by 0.85%, reaching 76.95% compared to 76.1%.
+These results demonstrate the scalability of our method to larger models and longer pretraining schedules.
+We emphasize that HVP is insensitive to the baseline hyperparameters and simply reusing the default ones consistently resulted in improvements in the reported magnitudes across all experiments.
 5.2 Transfer to Other Datasets and Tasks
-We now demonstrate the transferability of features learned with HVS. For all transfer experiments, we use our 100-epoch ImageNet pretrained SimSiam ResNet-50 and DINO ViT-S/16 models.
+We now report the transferability of features learned with HVP. For all experiments here, we use our 100-epoch ImageNet pretrained iBOT and DINO ViT-S/16 models, respectively.
 Method Arch. CIFAR10 CIFAR100 Flowers102 iNat 21 Food101
@@
 SimSiam RN50 82.60 95.50 54.20 77.20 34.27 56.40 32.50 60.30 65.70 83.90
-+ HVS RN50 84.40 96.10 57.10 78.20 38.37 58.90 33.90 60.90 67.10 84.70
-Improvement +1.80 +0.60 +2.90 +1.00 +4.10 +2.50 +1.40 +0.60 +1.40 +0.80
++ HVP RN50 84.40 96.10 57.10 78.20 38.37 58.90 33.90 60.90 67.10 84.70
+Impr. +1.80 +0.60 +2.90 +1.00 +4.10 +2.50 +1.40 +0.60 +1.40 +0.80
 DINO ViT-S/16 94.53 98.53 80.63 87.90 91.10 93.20 46.93 53.97 73.30 87.50
-+ HVS ViT-S/16 95.13 98.65 81.27 88.23 92.07 93.60 49.03 54.16 74.13 87.91
-Improvement +0.60 +0.12 +0.63 +0.33 +0.97 +0.40 +2.10 +0.19 +0.83 +0.41
-Table 2: Models trained on ImageNet with HVS compare favourably against models trained without it when transferred to other datasets through finetuning (F.T.) or the linear evaluation protocol (Lin.).
-Linear Evaluation and Finetuning
-In Table 2, we apply both the linear evaluation (Lin.) and finetuning (F.T.) protocols to our models across a diverse set of datasets consisting of CIFAR10 Krizhevsky (2009), CIFAR100, Flowers102Nilsback & Zisserman (2008), Food101Bossard et al.
-(2014), and iNaturalist 2021 (iNaturalist 2021 competition dataset, ).
-Our results show that the improvements achieved by sampling hard views we observed so far also transfer to other datasets.
-Method Arch. VOC Det. COCO Segm.
-SimSiam RN50 78.46 45.16
-+ HVS RN50 79.06 46.51
-Impr. +0.60 +1.35
-DINO RN50 80.86 50.35
-+ HVS RN50 80.51 50.39
-Impr. -0.36 +0.04
-Table 3: AP50 for object detection and instance segmentation.
-Object Detection and Instance Segmentation
-For object detection, we use the VOC07+12 (Everingham et al., 2010) dataset using Faster R-CNN (Ren et al., 2015).
-We study instance segmentation on the COCO (Lin et al., 2014) dataset using Mask R-CNN (He et al., 2022).
-Table 3, where we report the AP50 performance, shows that the features learned with HVS also mostly transfer favourably to these tasks and outperform the SimSiam baseline with a 100-ep.
++ HVP ViT-S/16 95.13 98.65 81.27 88.23 92.07 93.60 49.03 54.16 74.13 87.91
+Impr. +0.60 +0.12 +0.63 +0.33 +0.97 +0.40 +2.10 +0.19 +0.83 +0.41
+Table 2: HVP compares favorably against models trained without it when fine-tuned (F.T.) to or linearly evaluated (Lin.) on other datasets (averaged over 3 seeds; 100-ep. preraining).
+5.2.1 Linear Evaluation and Finetuning
+Method Arch. OD IS
+100 300 100 300
+iBOT ViT-S/16 66.13 66.80 63.10 63.63
++ HVP ViT-S/16 66.50 67.13 63.50 64.23
+Impr. +0.37 +0.33 +0.40 +0.60
+DINO ViT-S/16 65.90 66.60 62.83 63.63
++ HVP ViT-S/16 66.37 67.00 63.37 64.00
+Impr. +0.47 +0.40 +0.53 +0.37
+Table 3: Object Detection (OD) and Instance Segmentation (IS) AP50 performance on COCO for 100/300 epoch pretraining.
+In Table 2, we apply both the linear evaluation (Lin.) and finetuning (F.T.) protocols to our models across a diverse set of datasets consisting of CIFAR10 (Krizhevsky, 2009), CIFAR100, Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), and iNaturalist 2021 (iNaturalist 2021 competition dataset, ).
+Our results show that the improvements achieved by sampling hard views that we observed so far also transfer to other datasets.
+5.2.2 Object Detection and Instance Segmentation
+For object detection and instance segmentation, we use the COCO (Lin et al., 2014) dataset with Cascade Mask R-CNN (Cai & Vasconcelos, 2019; He et al., 2017).
+Table 3, where we report the AP50 performance, shows that the features learned with HVP also transfer favorably to these tasks and outperform the iBOT and DINO baseline with a 100-ep.
+and 300-ep.
 pretraining.
-However, in case of DINO, we observe a slight performance gap on the detection task, which is the only instance in the entire paper where HVS did not outperform the baseline.
-This result warrants further investigation, as it was conducted with a single seed (due to limited compute resources).
-More details and performance results on this task are provided in Appendix F.1.
-4.3 Time Complexity of HVS
-Method Arch. Lin.
-DINO ViT-S/16 74.06
-+ HVS ViT-S/16 74.67
-Impr. +0.61
-SimSiam RN50 68.46
-+ HVS RN50 68.98
-Impr. +0.52
-SimCLR RN50 64.31
-+ HVS RN50 65.33
-Impr. +1.02
-Table 4: Accounting for computational overhead with 50% resolution (100 epochs).
-The additional forward passes that HVS introduces for the selection phase increase the time complexity of the individual baseline methods.
-Several possible approaches exist to mitigate this overhead, one of which is to reduce the image resolution for the selection step (while still using the original resolution for learning).
-We measured the overhead factors under varying image resolutions for SimSiam and DINO and report them in Table 14 in the appendix.
-Given the 20% (SimSiam) to 30% (DINO) faster training times per batch when using a 50% image reduction, we evaluated this approach and observed that final performances are retained across all tested methods.
-Table 4 provides insights into the HVS’ performance when computational overhead is accounted for in the 50% reduction case.
-Overall, the results validate the effectiveness of HVS in reducing computational demands without compromising on performance.
-We refer to Section H in the appendix for a discussion on further optimizing HVS.
-5 Empirical Analysis of Hard View Selection
-In this section, we discuss studies designed to shed light on the mechanisms underlying HVS.
-We address the following questions: 1) “Which can be observed that underlie the hard view selection?” 2) “Can a “manual” augmentation policy be inferred from these patterns?”, and 3) “What are the effects on empowering the adversary?”.
-For all experiments conducted here, we use our 100-epoch, ImageNet-pretrained SimSiam+HVS models with four sampled views.
-5.1 Q1: Which Patterns Underlying the Hard View Selection Can be Observed?
-When visually studying examples and the views selected by HVS in Figures 2 and 5 (in the appendix), we notice that both geometric and appearance characteristics seem to be exploited, for instance, see the brightness difference between the views of the first two rows in Fig.
-2.
+More details and performance results on this task are provided in Appendix H.1.
+6 Empirical Analysis of HVP
+In this section, we discuss studies designed to shed light on the mechanisms underlying HVP.
+We address the following questions: 1) “Which pattern can be observed that underlies the hard view selection?” and 2) “What are the effects on empowering the adversary?”.
+For all experiments conducted here, we use our 100-epoch, ImageNet-pretrained SimSiam+HVP models with four sampled views.
+In Appendix F.1, we also analyze whether we can infer a “manual" augmentation policy from the following observed patterns.
+6.1 Q1: Which Patterns Can be Observed with HVP?
+When visually studying examples and the views selected by HVP in Figures 1(b) and 5 (in the appendix), we notice that both geometric and appearance characteristics seem to be exploited, for instance, see the brightness difference between the views of the first two rows in Fig.
+1(b).
 We also see a generally higher training loss (Fig.
 6 in the appendix) indicative of an increased task difficulty.
-Logging Augmentation Data
-To assess these observations, we logged relevant hyperparameter data during SimSiam training with HVS.
-The logs include for each view the sampled geometric and appearance parameters from the data augmentation operations (such as the height/width of the crops or the brightness; see Section D in the appendix for more details), as well as the loss and whether the view was selected.
-As evaluated metrics we chose the Intersection over Union (IoU), Relative Distance (normalized distance of the center points views), color distortion distance (the Euclidean distance between all four color distortion parameters), and the individual color distortion parameters.
-Figure 3: Left: In over 40% of the cases, the adversarially selected view pair is also the one with the lowest IoU throughout SimSiam+HVS pretraining.
-We attribute the spike in the early phase to the random initialization of the embedding.
-Right: When using HVS (blue), a shift to smaller IoU values is visible when comparing to standard pretraining (orange).
+Figure 2: Left: In over 40% of the cases, the adversarially selected view pair has also the lowest Intersection over Union throughout SimSiam+HVP pretraining.
+We attribute the early spike to the random initialization of the embedding.
+Right: HVP (blue) shows a shift to smaller IoU values over standard pretraining (orange).
 Both results are based on 3 seeds.
-Importance of Augmentation Metrics
+6.1.1 Logging Augmentation Data
+To assess these observations, we logged relevant hyperparameter data during SimSiam training with HVP.
+The logs include for each view the sampled geometric and appearance parameters from the data augmentation operations (such as the height/width of the crops or the brightness; see Section E in the appendix for more details), as well as the loss and whether the view was selected.
+As evaluated metrics, we chose the Intersection over Union (IoU), Relative Distance (normalized distance of the center points views), color distortion distance (the Euclidean distance between all four color distortion parameters), and the individual color distortion parameters.
+6.1.2 Importance of Augmentation Metrics
 Given 300k such samples, we then used fANOVA (Hutter et al., 2014) to determine how predictive these metrics are.
-This resulted in the metric with the highest predictive capacity on the loss to be the IoU, explaining 15% of the variance in performance, followed by brightness with 5% (for more details see Fig.
+This resulted in the metric with the highest predictive capacity on the loss being the IoU, explaining 15% of the variance in performance, followed by brightness with 5% (for more details see Fig.
 8 in the appendix).
-The importance of IoU in HVS is further underpinned by the following observation: the fraction of view pairs selected by HVS, which are also the pairs with the lowest IoU among all six pairs (N=4), is over 40% (random: ${\sim}16.7\%$) during training.
-Moreover, when using HVS, a shift to smaller IoU values can be observed when comparing against standard pretraining.
-We visualize both findings in Fig.
-3.
-Taking a Closer Look at the Intersection over Union
+The importance of IoU in HVP is further underpinned by the following observation: the fraction of view pairs selected by HVP, which are also the pairs with the lowest IoU among all six pairs (N=4), is over 40% (random: ${\sim}16.7\%$) during training.
+Moreover, when using HVP, a shift to smaller IoU values can be observed when comparing against standard pretraining (see Fig.
+2).
+Figure 3: Left: The average IoU of view pairs selected by SimSiam+HVP (blue) compared against the default SimSiam training (green).
+Right: Using static color augmentation for all pairs before the selection increases the dependency on the IoU.
+6.1.3 Taking a Closer Look at the Intersection over Union
 We also examined the IoU value over the course of training in Fig.
 3 (left).
-An observable pattern is that the IoU value with HVS (Fig.
-4 (left) in blue) is smaller and varies more when compared against training without HVS (green).
-We believe this is due to the sample-wise and stateful nature of the adversarial selection as HVS chooses different IoU values between varying samples and model states.
+An observable pattern is that the IoU value with HVP (Fig.
+3 (left) in green) is smaller and varies more when compared against training without HVP (blue).
+We believe this is due to the sample-wise and stateful nature of the adversarial selection as HVP chooses different IoU values between varying samples and model states.
 Lastly, we assessed the effect of the color augmentation on the pair selection.
 For this study, we sampled one set of color augmentations (as opposed to one for each view) per iteration and applied it to all views.
-Only after identifying the hardest pair, we apply sampled data augmentations to each view.
+We apply sampled data augmentations to each view only after identifying the hardest pair.
 As we show in Fig.
 3 (right), the fraction of selected pairs that are also the hardest pairs slightly increases in this case.
-One possible explanation for this is that it reflects the non-negligible role of color variation between views (as shown previously with the importance analysis), where HVS is given less leverage to increase hardness through a static appearance and instead, depends more on leveraging the IoU.
-Another key observation is that HVS often chooses view pairs that incorporate zooming in and out or an increased distance between the views (see last row of Fig.
-2).
-We believe the former relates to the analogy of local-to-global view correspondence described in SimCLR and the DINO papers.
-5.2 Q2: Can a Manual Augmentation Policy be Inferred?
+One possible explanation for this is that it reflects the non-negligible role of color variation between views (as shown previously with the importance analysis), where HVP is given less leverage to increase hardness through a static appearance and instead, depends more on leveraging the IoU.
+Another key observation is that HVP often chooses view pairs that incorporate zooming in and out or an increased distance between the views (see last row of Fig.
+1(b)).
+6.2 Q2: What are the Effects of Strong Adversaries?
+It is well known that adversarial learning can suffer from algorithmic instability (Xing et al., 2021), e.g.
+by giving an adversary too much capacity.
+Here, we further explore the space of adversarial capacity for pretraining with hard views by adapting and varying HVP hyperparameters in order to gain a better understanding of their impact and robustness on discriminative learning.
+Additionally, we report further results on learning an adversary in Appendix G.1.
+6.2.1 Robustness to Augmentation Hyperparameters
+Figure 4: With HVP, SimSiam appears more robust to augmentation hyperparameter variation.
+We show this for RandomResizedCrop (left) and ColorJitter (right).
+For RRC, the values indicate the lower value of the sampling range and for CJ the intensity of the color cues.
+Results averaged over two seeds and SimSiam defaults are RRC=0.2 and CJ=0.5.
+In assessing the robustness of HVP to augmentation hyperparameters, we conducted an analysis focusing on two primary augmentation operations: RandomResizedCrop (RRC) and ColorJitter (CJ).
+Our findings depicted in Fig.
+4 suggest that HVP enhances the robustness of SSL methods, like SimSiam, against variations in these hyperparameters.
+Note, when varying one operation, either RRC or CJ, we maintained the other operation at its default configuration.
+In both settings, we observe less performance degradation with extreme augmentation values and overall smaller degradation rates for HVP.
+We believe that this robustness stems from hard view pretraining which inadvertently equips the model to handle stronger augmentations.
+6.2.2 Increasing the Number of Views
+In our initial experiments, we explored variations in the number of sampled views $N$ with SimSiam and HVP.
+As can be seen in Fig.
+7 in the appendix, while $N=8$ views still outperform the baseline in terms of linear evaluation accuracy, it is slightly worse than using $N=4$ views (-0.05% for 100 epochs pretraining on linear evaluation).
+We interpret this result as the existence of a “sweet spot" in setting the number of views, where, in the limit, a higher number of views corresponds to approximating a powerful adversarial learner, capable of choosing very hard and unfavorable learning tasks that lead to model collapse and performance deterioration.
+We experimented with such an adversarial learner and report results in Appendix G.1.
+7 Conclusion
+We presented HVP, a new data augmentation and learning strategy for Self-Supervised Learning designed to challenge pretrained models with harder samples.
+This straightforward method allows pushing the effectiveness of the traditional random view generation in SSL.
+When combined with methods like DINO, SimSiam, iBOT, and SimCLR, HVP consistently showcased improvements of 1% on average in linear evaluation and a diverse set of transfer tasks.
+HVP achieved a new state-of-the-art result of 78.8% linear evaluation accuracy on DINO ViT-B/16, a 0.6% improvement over the previous baseline.
+This illustrates the scalability and effectiveness of our approach for larger pretraining settings.
+With growing models, there is an increasing demand for more data to effectively train them.
+Synthetic data generation offers a viable solution by enhancing the quantity and diversity of training data.
+Data augmentation techniques like HVP play a crucial role in this process by creating challenging views, which can serve as synthetic data.
+All in all, HVP holds promise for scenarios where one seeks to push absolute performance or explore making models less sensitive to hyperparameters, thereby strengthening them for various downstream applications.
+8 Acknowledgements
+Frank Hutter acknowledges the financial support of the Hector Foundation.
+We also acknowledge funding by the European Union (via ERC Consolidator Grant DeepLearning 2.0, grant no.
+101045765).
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council.
+Neither the European Union nor the granting authority can be held responsible for them.
+Moreover, we acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828, by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG, as well as the Gauss Center for Supercomputing eV (www.gauss-centre.eu) for funding this project by providing computing time on the GCS supercomputer JUWELS at Jülich Supercomputing Center (JSC).
+We also acknowledge Robert Bosch GmbH for financial support and the computing time made available on the high-performance computer NHR@KIT Compute Cluster at the NHR Center NHR@KIT.
+Appendix
+Appendix A Reproducibility Statement
+We provide complete code, environment installation instructions, hyperparameter settings and model checkpoints here: https://github.com/automl/hvp.
+For transparency, we outline all hyperparameters, data splits, and evaluation protocols in detail in Section I.
+Most experiments were run across multiple seeds, and we report average results to account for variability.
+Information regarding the required computational resources is discussed in Section J below.
+Appendix B Examples Sampled by HVP
+Figure 5: We depict row-wise ten example images from the ImageNet train set along with their sampled views with a finished, 100-epoch trained SimSiam ResNet50 model.
+Left: original image with the overlaid randomly sampled crops (colored dashed rectangles).
+Right: All views after applying resizing and appearance augmentations.
+The pair that is selected adversarially by HVP is highlighted in solid lines, eg.
+View 1 and View 4 in the first row.
+Appendix C Training Loss
+Figure 6: The training loss over 100 epochs.
+Comparing the DINO vanilla method with DINO + HVP.
+The spike and drop in the loss curve of DINO is caused by freezing the last layer in the first epoch which was proposed by the authors as a strategy to enhance downstream performance.
+For HVP we can only see a drop and no spike.
+We believe this is because HVP exposes the model to hard views from the beginning of training (i.e.
+the loss is immediately maximized).
+Appendix D Effect of More Views on Linear Evaluation Performance
+Figure 7: Setting the number of views too high can result in performance deterioration.
+This shows that diminishing returns exist, likely because the adversary becomes too strong, resulting in a too hard learning task.
+Appendix E Assessing the Importance of Metrics with fANOVA
+To assess the importance of various metrics on the training loss, we apply fANOVA Hutter et al.
+(2014) on data that we logged during training with HVP.
+We used 300k samples that contain the following sampled parameters from the geometric and appearance data augmentation operations for each view: all random resized crop parameters (height and width of the original image, coordinates of crop corners and height and width of the crop), all Colorjitter (color distortion) strengths (brightness, contrast, saturation, hue), grayscale on/off, Gaussian blurring on/off, horizontal flip on/off, loss, and if the crop was selected or not.
+The metrics we chose are Intersection over Union (IoU), Relative Distance (sample-wise normalized distance of the center points of crop pairs), color distortion distance (the Euclidean distance between all four color distortion operation parameters, i.e.
+brightness, contrast, saturation, hue), and the individual color distortion parameters of the Colorjitter operation.
+As can be seen in Fig.
+8, the metric with the highest predictive capacity on the loss is the IoU with an importance of 15.2% followed by brightness with 5.1%.
+The relative distance has an importance of 3.3%, the Colorjitter distance 2.3%, the contrast 1.6%, the saturation 1.4%, hue 0.6%, and all parameters jointly 1.7%.
+Figure 8: Application of fANOVA Hutter et al. (2014) on logged training data to determine metrics with high predictive capacity on the training loss.
+Appendix F Additional Empirical Analysis
+F.1 Can a Manual Augm. Policy be Derived?
 IoU Policy Type SimSiam DINO
@@
 B+range(1.0-0.1) +0.07 -
-Table 5: Top-1 lin. eval. accuracies for the manual IoU policy (averaged over 2 seeds).
-Since harder pretraining tasks seem beneficial according to observations made in Q1, a natural question arises: is it possible to mimic the adversarial selection with a manually scripted augmentation policy?
-Such a policy would replace HVS and lower the computational cost .
-Since the IoU plays an important role, below, we study several possible ways for constructing a simple manual augmentation policy based on IoU.
-Deriving an Augmentation Policy
+Table 4: Top-1 lin. eval. accuracies for the manual IoU policy.
+Since harder pretraining tasks seem beneficial according to observations made in Q1, a natural question arises: can we mimic the adversarial selection with a manually scripted augmentation policy?
+Such a policy would replace HVP and lower the computational cost.
+Since the IoU plays an important role, below, we study several ways to construct a simple manual augmentation policy based on IoU.
+F.1.1 Deriving an Augmentation Policy
 We implemented the following rejection sampling algorithm in the augmentation pipeline: we linearly approximate the IoU values from Fig.
@@
 We varied different hyperparameters, e.g., IoU start/end ranges, inverse schedules, and alternating between the IoU schedule and the standard augmentation .
-Training both SimSiam and DINO models for 100 epochs yielded performance decreases or insignificant improvements (see Table 5).
-These results indicate that developing a manual policy based on metrics in pixel space such as the IoU is non-trivial.
+Training both SimSiam and DINO models for 100 epochs yielded performance drops or insignificant improvements (see Table 4).
+These results indicate that using a manual policy based on metrics in pixel space such as the IoU is non-trivial.
 Additionally, these results show that transferring such a policy from SimSiam to DINO does not work well, possibly due to additional variations in the augmentation pipeline such as multi-crop.
-In contrast, we believe that HVS is effective and transfers well since it 1) operates on a similarity level of latent embeddings that may be decorrelated from the pixel space and 2) has access to the current model state.
-Figure 4: Left: The average IoU of view pairs selected by SimSiam+HVS (blue) compared against the default SimSiam training (green).
-Right: Using static color augmentation for all pairs before the selection increases the dependency on the IoU.
-Assessing the Difficulty of Predicting the Pair Selected by HVS
-We further validated the previous result and assess an upper limit on the performance for predicting the hardest pair.
-By fitting a gradient boosting classification tree (Friedman, 2001), we predict the selected view pair conditioned on all SimSiam hyperparameter log data from Q1 except for the flag that indicates whether a view was selected.
+In contrast, we believe that HVP is effective and transfers well since it 1) operates on a similarity level of latent embeddings that may be decorrelated from the pixel space and 2) has access to the current model state.
+F.1.2 Assessing the Difficulty of Predicting the Pair Selected by HVP
+We further validate the previous result and assess an upper limit on the performance for predicting the hardest pair.
+By fitting a gradient boosting classification tree Friedman (2001), we predict the selected view pair conditioned on all SimSiam hyperparameter log data from Q1 except for the flag that indicates whether a view was selected.
 As training and test data, we used the logs from two seeds (each 300k samples) and the logs from a third seed, respectively.
@@
 However, the resulting average test performance in all scenarios never exceeded 40%, indicating that it is indeed challenging to predict the hardness of views based on parameter-level data.
-This further supports our hypothesis that deriving a policy for controlling and increasing hardness based on geometric and appearance parameters is non-trivial and that such a policy must function on a per-sample basis and have access to the current model state as in HVS.
-5.3 Q3: What are the Effects of Empowering the Adversary?
-It is well known that adversarial learning can suffer from algorithmic instability (Xing et al., 2021), e.g.
-by giving an adversary too much capacity.
-We now further explore the space of adversarial capacity for selecting hard views to gain a better understanding of its impact on contrastive learning.
-Optimizing an Adversarial Learner for the View Generation
-In this experiment, we briefly explore adversarially learning a network to output the transformation matrix for view generation.
-We optimize a Spatial Transformer Network (STN) (Jaderberg et al., 2015) jointly with the DINO objective and a ViT-tiny.
-Using the inverse of the DINO loss, we train the STN to output the geometric transformation matrices (each 6 parameters) for generating views .
-We employ and test various transformation types (e.g., rotation, translation, etc.) that influence the capacity of the adversary.
-We discuss this experiment in greater detail in Appendix E and briefly summarize our findings here: applying an adversary in this way is very challenging due to the sensitivity of CL on the data augmentation and the dynamics of adversarial learning, often leading to either minor improvements or model collapses as we have observed when employing a too strong manual IoU policy in Q1.
-Increasing the Number of Views
-In our initial experiments, we explored variations in the number of sampled views $N$ with SimSiam and HVS.
-As can be seen in Fig.
-7 in the appendix, while $N=8$ views still outperform the baseline in terms of linear evaluation accuracy, it is slightly worse than using $N=4$ views (-0.05% for 100 epochs and -0.14% for 200 epochs pretraining on linear evaluation).
-We interpret this result as the existence of a “sweet spot” in setting the number of views, where, in the limit, a higher number of views corresponds to approximating the adversarial learner described above.
-This would allow HVS to select progressively harder views that quickly results in very hard tasks and consequently in model collapse and performance deterioration.
-[^1] For completeness, we also experimented with a cooperative HVS (see Appendix F.2 for details).
-6 Conclusion
-In this study, we presented HVS, a new data augmentation and learning strategy for Self-Supervised Learning designed to challenge pretrained models with more demanding samples.
-This method, while straightforward in its design, has proven to be a powerful tool, pushing the boundaries of the traditional random view generation in SSL.
-When combined with methods like DINO, SimSiam, iBOT, and SimCLR, HVS consistently showcased improvements in linear evaluation and a diverse set of transfer tasks.
-As the landscape of self-supervised learning continues to evolve, methods like HVS underscore the potential for further advancements in the field.
-By continuously challenging models with harder samples and refining augmentation techniques, we anticipate even greater strides in model performance and generalization in future research endeavors.
-While HVS introduces computational overhead due to its additional forward passes, our findings demonstrate that this overhead is significantly mitigated when using small resolution images in HVS without impeding its performance.
-This suggests that HVS is likely to continue to outperform baselines regardless of the training length.
-Appendix A Examples Sampled by HVS
-Figure 5: We depict row-wise ten example images from the ImageNet train set along with their sampled views with a finished, 100-epoch trained SimSiam ResNet50 model.
-Left: original image with the overlaid randomly sampled crops (colored dashed rectangles).
-Right: All views after applying resizing and appearance augmentations.
-The pair that is selected adversarially by HVS is highlighted in solid lines, eg.
-View 1 and View 4 in the first row.
-Appendix B Training Loss
-Figure 6: The training loss over 100 epochs.
-Comparing the DINO vanilla method with DINO + HVS.
-The spike and drop in the loss curve of DINO is caused by freezing the last layer in the first epoch which was proposed by the authors as a strategy to enhance downstream performance.
-For HVS we can only see a drop and no spike.
-We believe this is because HVS exposes the model to hard views from the beginning of training (i.e.
-the loss is immediately maximized).
-Appendix C Effect of More Views on Linear Evaluation Performance
-Figure 7: Setting the number of views too high can result in performance deterioration.
-This shows that diminishing returns exist, likely because the adversary becomes too strong, resulting in a too hard learning task.
-Appendix D Assessing the Importance of Metrics with fANOVA
-To assess the importance of various metrics on the training loss, we apply fANOVA Hutter et al.
-(2014) on data that we logged during training with HVS.
-We used 300k samples that contain the following sampled parameters from the geometric and appearance data augmentation operations for each view: all random resized crop parameters (height and width of the original image, coordinates of crop corners and height and width of the crop), all Colorjitter (color distortion) strengths (brightness, contrast, saturation, hue), grayscale on/off, Gaussian blurring on/off, horizontal flip on/off, loss, and if the crop was selected or not.
-The metrics we chose are Intersection over Union (IoU), Relative Distance (sample-wise normalized distance of the center points of crop pairs), color distortion distance (the Euclidean distance between all four color distortion operation parameters, i.e.
-brightness, contrast, saturation, hue), and the individual color distortion parameters of the Colorjitter operation.
-As can be seen in Fig.
-8, the metric with the highest predictive capacity on the loss is the IoU with an importance of 15.2% followed by brightness with 5.1%.
-The relative distance has an importance of 3.3%, the colorjitter distance 2.3%, the contrast 1.6%, the saturation 1.4%, hue 0.6%, and all parameters jointly 1.7%.
-Figure 8: Application of fANOVA Hutter et al. (2014) on logged training data to determine metrics with high predictive capacity on the train loss.
-Appendix E Adversarial Learner for the View Generation
-We use Spatial Transformer Network (STN) (Jaderberg et al., 2015) to allow generating views by producing 6D transformation matrices (allowing translating, rotating, shearing, scaling, affine transformations and combinations thereof) in a differentiable way since most common augmentations are not off-the-shelf differentiable.
-As described, we train it alongside the actual pretrained network using the same (inverted) objective.
-For our experiments we use DINO with multi-crop, i.e.
+This further supports our hypothesis that deriving a policy for controlling and increasing hardness based on geometric and appearance parameters is non-trivial and that such a policy must function on a per-sample basis and have access to the current model state as in HVP.
+Appendix G Learning View Generation
+G.1 Adversarial Learner
+In this experiment, we explore adversarially learning a network to output the transformation matrix for view generation.
+We use Spatial Transformer Network (STN) Jaderberg et al.
+(2015) to allow generating views by producing 6D transformation matrices (allowing translating, rotating, shearing, scaling, affine transformations, and combinations thereof) in a differentiable way since most common augmentations are not off-the-shelf differentiable.
+We optimize the STN jointly with the DINO objective and a ViT-tiny.
+We train it alongside the actual pretrained network using the same (inverted) objective.
+For our experiments, we use DINO with multi-crop, i.e.
 2 global and 8 local heads.
-As STN we use a small CNN with a linear layer for outputting the 10*6D transformation matrices.
+As STN we use a small CNN followed by a linear layer for outputting the 10*6D transformation matrices.
 In this scenario, we use a ViT-tiny/16 with a 300 epoch pretraining on CIFAR10 with a batch size of 256.
@@
 Figure 9: Illustration of adversarial learning with a Spatial Transformer Network (STN) jointly with contrastive learning (here: DINO).
-As mentioned previously, the STN, without using auxiliary losses, starts zooming in and generating single-color views.
+The STN, without using auxiliary losses, starts zooming in and generating single-color views.
 To counteract this behavior, we experimented with different penalties on the transformation matrices produced by the STN.
@@
 This is on the one hand due to the penalties not being fully able to restrict the output of the STN.
-On the other, the task of extracting useful information from two differently rotated crops is even harder, and learning spatial invariance becomes to too challenging.
+On the other, the task of extracting useful information from two differently rotated crops is even harder, and learning spatial invariance becomes too challenging.
 All in all, we experienced two modes: either the STN is too restricted, leading to static output (i.e.
@@
 See Fig.
-10 for an example on the former behavior.
+10 for an example of the former behavior.
 Figure 10: Example for static output behavior of the STN.
@@
 Table 5: Linear evaluation and finetuning classification performance on CIFAR10. Top-1 accuracy on the validation set of CIFAR10 for our best results reported with different STN transformation modes.
-Appendix F Additional Results
-F.1 Object Detection and Instance Segmentation
-For our object detection and instance segmentation analysis experiment, we report additional results in Table F.1 based on the AP metric.
-We use the C4 backbone variant (Wu et al., 2019) and finetuning with the 1x schedule.
-Method Arch. VOC07+12 Object Det. COCO Instance Segm.
-$\mathbf{AP_{all}}$ $\mathbf{AP_{50}}$ $\mathbf{AP_{75}}$ $\mathbf{AP_{mask}}$ $\mathbf{AP_{mask50}}$ $\mathbf{AP_{mask75}}$
-SimSiam RN50 52.05 78.46 56.88 27.16 45.16 28.48
-+ HVS RN50 52.91 79.06 57.55 27.93 46.51 29.36
-Improvement +0.86 +0.60 +0.67 +0.77 +1.35 +0.88
-DINO RN50 53.66 80.86 58.60 30.25 50.35 31.78
-+ HVS RN50 52.86 80.51 58.21 30.02 50.39 31.52
-Improvement +0.80 -0.36 -0.39 -0.23 +0.04 -0.27
-Table 7:
-F.2 Easy View Selection
+G.2 Cooperative Learner
 To investigate the effect of a cooperative, i.e.
@@
 This result is in line with previous findings that highlight the importance of strong augmentations in CL.
-Appendix G Hyperparameters
-G.1 Evaluations on ImageNet
-G.1.1 DINO
-For DINO, we report the ViT pretraining hyperparameters in Table 8.
-For ResNet-50, we use the same hyperparameters.
-Note, for HVS we limit the total number of comparisons to 128 across all heads.
-Linear evaluation is executed for 100 epochs and we use a learning rate of $0.001$, SGD optimizer (AdamW (Loshchilov & Hutter, 2019) during pretraining), a batch size of 1024, a momentum of $0.9$, and no weight decay.
+Appendix H Additional Results
+H.1 Object Detection and Instance Segmentation
+Here we provide more detailed results on object detection and instance segmentation, shown in Tab.
+6 and Tab.
+7 respectively.
+We followed iBOT’s default configurations which employed Cascade Mask R-CNN as the task layer.
+Method Arch. Object Detection Instance Segmentation
+$\mathbf{AP^{b}}$ $\mathbf{AP^{b}_{50}}$ $\mathbf{AP^{b}_{75}}$ $\mathbf{AP^{m}}$ $\mathbf{AP^{m}_{50}}$ $\mathbf{AP^{m}_{75}}$
+iBOT ViT-S/16 47.00 66.13 50.63 40.67 63.10 43.37
++ HVP ViT-S/16 47.27 66.50 50.90 40.90 63.50 43.83
+Improvement +0.27 +0.37 +0.27 +0.23 +0.40 +0.47
+DINO ViT-S/16 46.50 65.90 50.30 40.43 62.83 43.27
++ HVP ViT-S/16 47.07 66.37 50.63 40.80 63.37 43.87
+Improvement +0.57 +0.47 +0.33 +0.37 +0.53 +0.60
+Table 6: Additional object detection and instance segmentation results on the COCO dataset. The ViT-S/16 models were pretrained for 100 epochs.
+Method Arch. Object Detection Instance Segmentation
+$\mathbf{AP^{b}}$ $\mathbf{AP^{b}_{50}}$ $\mathbf{AP^{b}_{75}}$ $\mathbf{AP^{m}}$ $\mathbf{AP^{m}_{50}}$ $\mathbf{AP^{m}_{75}}$
+iBOT ViT-S/16 47.60 66.80 51.33 41.10 63.63 44.07
++ HVP ViT-S/16 48.03 67.13 51.73 41.50 64.23 44.40
+Improvement +0.43 +0.33 +0.40 +0.40 +0.60 +0.33
+DINO ViT-S/16 47.27 66.60 51.00 41.00 63.63 44.03
++ HVP ViT-S/16 47.50 67.00 51.33 41.17 64.00 44.13
+Improvement +0.23 +0.40 +0.33 +0.17 +0.37 +0.10
+Table 7: Additional object detection and instance segmentation results on the COCO dataset. The ViT-S/16 models were pretrained for 300 epochs.
+Appendix I Hyperparameters
+I.1 Evaluations on ImageNet
+I.1.1 DINO
+For DINO, we report the ViT pretraining hyperparameters in Table 8 (ViT-S) and Table 9 (ViT-B) which are the original ones as reported by the authors.
+Note, for HVP we limit the total number of comparisons to 128 across all heads.
+Linear evaluation is executed for 100 epochs and we use a learning rate of $0.00075$, SGD optimizer (AdamW Loshchilov & Hutter (2019) during pretraining), a batch size of 1024, a momentum of $0.9$, and no weight decay (as reported by the authors).
 Hyperparameter Value Hyperparameter Value
-architecture vit_small epochs: 100
-img_size 224 warmup_epochs: 10
-patch_size 16 freeze_last_layer: 1
-out_dim 65536 lr: 0.0005
-norm_last_layer true min_lr: 1.0e-06
-momentum_teacher 0.996 optimizer: AdamW
-use_bn_in_head false weight_decay: 0.04
-teacher_temp 0.04 weight_decay_end: 0.4
-warmup_teacher_temp 0.04 global_crops_scale: 0.4, 1.0
-warmup_teacher_temp_epochs 0 global_crops_size: 224
-fp16 true local_crops_number: 8
-batch_size 512 local_crops_scale 0.05, 0.4
-clip_grad 3.0 local_crops_size: 96
-drop_path_rate 0.1
-Table 8: Pretraining ImageNet hyperparameters for the runs with DINO. For 300 epochs, we use a batch size of 1024.
-G.1.2 SimSiam
+architecture vit-small epochs: 100
+img-size 224 warmup-epochs: 10
+patch-size 16 freeze-last-layer: 1
+out-dim 65536 lr: 0.0005
+norm-last-layer true min-lr: 1.0e-06
+momentum-teacher 0.996 optimizer: AdamW
+use-bn-in-head false weight-decay: 0.04
+teacher-temp 0.04 weight-decay-end: 0.4
+warmup-teacher-temp 0.04 global-crops-scale: 0.4, 1.0
+warmup-teacher-temp-epochs 0 global-crops-size: 224
+fp16 true local-crops-number: 8
+batch-size 512 local-crops-scale 0.05, 0.4
+clip-grad 3.0 local-crops-size: 96
+drop-path-rate 0.1
+Table 8: Pretraining ImageNet hyperparameters for the runs with DINO ViT-S/16. For 300 epochs, we use a batch size of 1024.
+Hyperparameter Value Hyperparameter Value
+architecture vit-base epochs: 400
+img-size 224 warmup-epochs: 10
+patch-size 16 freeze-last-layer: 3
+out-dim 65536 lr: 0.00075
+norm-last-layer true min-lr: 2.0e-06
+momentum-teacher 0.996 optimizer: AdamW
+use-bn-in-head false weight-decay: 0.04
+teacher-temp 0.07 weight-decay-end: 0.4
+warmup-teacher-temp 0.04 global-crops-scale: 0.25, 1.0
+warmup-teacher-temp-epochs 50 global-crops-size: 224
+fp16 false local-crops-number: 10
+batch-size 1024 local-crops-scale: 0.05, 0.25
+clip-grad 0.3 local-crops-size: 96
+drop-path-rate 0.1
+Table 9: Pretraining ImageNet hyperparameters for the runs with DINO ViT-B/16.
+I.1.2 SimSiam
 In Table 10, we report the ResNet-50 pretraining hyperparameters.
-Linear evaluation is executed for 90 epochs (as reported by the SimSiam authors) and we use a learning rate of 0.1, LARS optimizer (You et al., 2017), a batch size of 4096, and no weight decay.
+Linear evaluation is executed for 90 epochs (as reported by the SimSiam authors) and we use a learning rate of 0.1, LARS optimizer You et al.
+(2017), a batch size of 4096, and no weight decay.
 Hyperparameter Value
 architecture resnet50
-batch_size 512
-blur_prob 0.5
-crops_scale 0.2, 1.0
-crop_size 224
-feature_dimension 2048
+batch-size 512
+blur-prob 0.5
+crops-scale 0.2, 1.0
+crop-size 224
+feature-dimension 2048
 epochs 100
-fix_pred_lr true
+fix-pred-lr true
 lr 0.05
 momentum 0.9
-predictor_dimension 512
-weight_decay 0.0001
+predictor-dimension 512
+weight-decay 0.0001
 optimizer SGD
 Table 10: Pretraining ImageNet hyperparameters for the runs with SimSiam. For 300 epochs, we use a batch size of 1024.
-G.1.3 SimCLR
-We report the ResNet-50 pretraining hyperparameters for SimCLR in Table 10. Linear evaluation is executed for 90 epochs with a learning rate 0.1, SGD optimizer, batch size of 4096 and no weight decay.
+I.1.3 SimCLR
+We report the ResNet-50 pretraining hyperparameters for SimCLR in Table 11.
+Linear evaluation is executed for 90 epochs with a learning rate of 0.1, SGD optimizer, batch size of 4096, and no weight decay.
 Hyperparameter Value
 architecture resnet50
-proj_hidden_dim 2048
-out_dim 128
-use_bn_in_head true
-batch_size 4096
+proj-hidden-dim 2048
+out-dim 128
+use-bn-in-head true
+batch-size 4096
 optim LARS
 lr 0.3
-sqrt_lr false
+sqrt-lr false
 momentum 0.9
-weight_decay 1e-4
+weight-decay 1e-4
 epochs 100
-warmup_epochs 10
-zero_init_residual true
+warmup-epochs 10
+zero-init-residual true
 Table 11:
-G.2 Transfer to Other Datasets and Tasks
+I.2 Transfer to Other Datasets and Tasks
 For linear evaluation on the transfer datasets, we simply used the same hyperparameters for linear evaluation on ImageNet (DINO and SimSiam respectively).
@@
 lr 7.5e-6 7.5e-6 5e-5 5e-5 5e-5
-weight_decay 0.05 0.05 0.05 0.05 0.05
+weight-decay 0.05 0.05 0.05 0.05 0.05
 optimizer AdamW AdamW AdamW AdamW AdamW
 epochs 300 300 300 100 100
-batch_size 512 512 512 512 512
+batch-size 512 512 512 512 512
 Table 12: Finetuning hyperparameters for DINO ViT-S/16.
@@
 lr 7.5e-6 5e-6 5e-4 7e-5 5e-6
-weight_decay 0.05 0.05 0.05 0.05 0.05
+weight-decay 0.05 0.05 0.05 0.05 0.05
 optimizer AdamW AdamW AdamW AdamW AdamW
 epochs 300 300 300 100 100
-batch_size 512 512 512 512 512
+batch-size 512 512 512 512 512
 Table 13: Finetuning hyperparameters for SimSiam and ResNet-50.
-G.3 Object Detection and Instance Segmentation
-We have used the Detectron2 library (Wu et al., 2019) for object detection and instance segmentation.
-We followed the public codebase from MoCo (He et al., 2020) (like SimSiam) for all entries.
-Due to limited compute resources we changed the batch size.
-All parameters that differ from MoCo are reported in the Table 13.
-The pretrained models are finetuned end-to-end on the target datasets.
-All methods are based on 100-epoch pre-training on ImageNet.
-VOC07+12 Object Det. Coce Inst. Segm.
-Hyperparameter Value Value
-batch_size 8 8
-lr 0.01 0.01
-steps 48000 90000
-Table 13: Hyperparameters for VOC object detection and COCO instance segmentation.
-Appendix H Computational Overhead of HVS
-In Table 14, we compare the computational overhead factors across multiple baselines.
-The details of hardware and software used for this analysis are: NVIDIA RTX 3080, AMD R7 5800X, 32GB RAM, Ubuntu 22.04, PyTorch 2.0.1, CUDA 11.8.
-For DINO’s 2 global and 8 local views (default), applying HVS with nviews=2 sampled for each original view results in 4 global and 16 local views.
-Since considering all combinations would yield over 77k unique comparisons (4 over 2 times 16 over 8), to remain tractable, we limit the number of total comparisons to 128.
-Overall, reducing the image resolution can significantly reduce the time complexity of HVS.
-For instance, in the case of DINO ViT-S/16, the reduction leads to approximately 30% faster training times and as reported in the main paper, a 50% reduction allows retaining original performance.
-Res. SimSiam (RN50) DINO (ViT)
-100% x1.55 x2.15
-75% x1.45 x1.68
-50% x1.24 x1.56
-25% x1.16 x1.5
-Table 14: Time slowdown factors under different image resolutions for the selection in HVS.
-While technically there can be a memory overhead with HVS, with the number of sampled views chosen in this paper, the backward pass of the methods that compute gradients only for the selected view pair still consumes more memory than the selection part of HVS (even for 8 sampled views in SimSiam).
+I.3 Object Detection and Instance Segmentation
+Our experiments utilized Open MMLab’s detection library Chen et al. (2019) for object detection and instance segmentation on COCO Lin et al. (2014). We followed iBOT’s default configuration.
+Obj. Det. & Inst. Segm. on COCO
+Hyperparameter Value
+epochs 12
+batch-size 32
+lr 0.02
+Table 14: Hyperparameters object detection and instance segmentation on COCO.
+Appendix J Computational Overhead of HVP
+The additional forward passes that HVP introduces for the selection phase increase the time complexity of the individual baseline methods.
+Several possible approaches exist to mitigate this overhead, one of which is to alternate between the vanilla and HVP training step.
+We measured the overhead factors for different alternating frequencies (i.e., after how many training steps we use the hard views from HVP; we refer to this as step) for SimSiam, DINO, and iBOT which we report in Table 15.
+Below, we propose additional ways that may allow using hard views more efficiently.
+Step
+SimSiam
+DINO
+(ViT-S/16)
+DINO
+iBOT
+(ViT-S/16)
+1 (HVP) x1.64 x2.21 x2.01 x2.13
+2 x1.38 x1.61 x1.56 x1.57
+3 x1.32 x1.43 x1.42 x1.38
+4 x1.29 x1.34 x1.35 x1.29
+Table 15: Slowdown factors for HVP and the alternating training method, where step refers to the interval at which HVP is applied during training (i.e., step=1 refers to full HVP and step=3 indicates that HVP is approx.
+used 33% of the training time).
+Measurements are averages across 3 seeds.
+As an alternative to the alternating training, we also experimented with a 50% image resolution reduction for the selection phase but observed that the final performance was negatively affected by it or that baseline performance could not be improved.
+The details of hardware and software used for this analysis are: one single compute node with 8 NVIDIA RTX 2080 Ti, AMD EPYC 7502 (32-Core Processor), 512GB RAM, Ubuntu 22.04.3 LTS, PyTorch 2.0.1, CUDA 11.8.
+For DINO’s 2 global and 8 local views (default), applying HVP with nviews=2 sampled for each original view results in 4 global and 16 local views.
+Since considering all combinations would yield over 77k unique comparisons ($\binom{4}{2}\times\binom{16}{8}$), to remain tractable, we limit the number of total comparisons to 64.
+For training experiments that exceed the limit of 8x RTX 2080 Ti, we apply gradient accumulation.
+While technically there can be a memory overhead with HVP, with the number of sampled views chosen in this paper, the backward pass of the methods that compute gradients only for the selected view pair still consumes more memory than the selection part of HVP (even for 8 sampled views in SimSiam).
 Note, that selection and the backward computation are never executed at the same time but sequentially.
-We emphasize that the time overhead factors were measured without any optimization of HVS’ efficiency and, in our view, there are multiple ways to improve it.
-In our work, we opted for the easiest implementation possible to showcase that selecting harder views dependent on the model learning state can help boost performance in contrastive learning.
-Going forward with more compute-efficient HVS solutions, one could think of:
+We emphasize that further ways exist to optimize HVP’s efficiency which remain to be explored. For instance:
   • using embeddings of views from “earlier” layers in the networks or
@@
   • using one GPU just for creating embeddings and selecting the hardest views while the remaining GPUs are used for learning or
-  • switching between HVS and the standard pipepline in alternating fashion or
-  • bypassing forwarding of similar pairs and more
-Appendix I Hard View Selection Objectives
-I.1 SimCLR
-In this section, we are going to introduce the application of HVS with the SimCLR objective.
+  • other approaches to derive manual augmentation policies or
+  • bypassing forwarding of similar pairs.
+Appendix K Hard View Pretraining Objectives
+K.1 SimCLR
+In this section, we are going to introduce the application of HVP with the SimCLR objective.
 Assume a given set of images $\mathcal{D}$, an image augmentation distribution $\mathcal{T}$, a minibatch of $M$ images $\mathbf{x}=\left\{x_{i}\right\}_{i=1}^{M}$ sampled uniformly from $\mathcal{D}$, and two sets of randomly sampled image augmentations $A=\left\{t_{i}\sim\mathcal{T}\right\}_{i=1}^{M}$ and $B$ sampled from $\mathcal{T}$.
@@
 Intuitively, when optimizing $\theta$, embeddings of two augmented views of the same image are attracted to each other while embeddings of different images are pushed further away from each other.
-To further enhance the training process, we introduce a modification to the loss function where instead of having two sets of augmentations $A$ and $B$, we now have ”N” sets of augmentations, denoted as $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{N}\}$.
-Each set $A_{i}$ is sampled from the image augmentation distribution $\mathcal{T}$, and applied to each image in $\mathbf{x}$, resulting in ”N” augmented sets of views $\mathbf{x}^{A_{1}},\mathbf{x}^{A_{2}},\ldots,\mathbf{x}^{A_{N}}$.
+To further enhance the training process, we introduce a modification to the loss function where instead of having two sets of augmentations $A$ and $B$, we now have "N" sets of augmentations, denoted as $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{N}\}$.
+Each set $A_{i}$ is sampled from the image augmentation distribution $\mathcal{T}$, and applied to each image in $\mathbf{x}$, resulting in "N" augmented sets of views $\mathbf{x}^{A_{1}},\mathbf{x}^{A_{2}},\ldots,\mathbf{x}^{A_{N}}$.
 Similarly, we obtain $N$ sets of embeddings $\mathbf{z}^{A_{1}},\mathbf{z}^{A_{2}},\ldots,\mathbf{z}^{A_{N}}$ through the encoder and projector networks defined as:
@@
 In other words, the hardest pair is the one, that has the lowest similarity with another augmented view of itself and the lowest dissimilarity with all other examples.
-Appendix J Attention Maps
-In this section, we visualize attention maps as a way to qualitatively study the potential effects and differences of the HVS learning method and their resulting features.
-The provided attention maps are from a DINO ViT-S/16 100 model, and we contrast the attention for each HVS and the baseline.
-All input images are from the ImageNet-1k validation split.
-The color code used depicts strong attention in yellow and weak attention in green and attentions from the HVS model are shown in the top row and attentions from the baseline in the bottom row, respectively.
-To summrize this study, we do not see apparent strong differences between HVS and the baseline.
-What we occasionally notice is that the HVS models seem to capture the shape of some subtle/indistinct objects better (e.g.
-the creek/river in the valley in Fig.
-12, the lizard in Fig.
-13, or the speaker in Fig.
-14 or focus slightly more on the context and background (Fig.
-16, 17, and 18).
-Figure 11:
-Figure 12:
-Figure 13:
-Figure 14:
-Figure 15:
-Figure 16:
-Figure 17:
-Figure 18:
-Figure 19: